### v1

# Efficient Online Inference of Vision Transformers by Training-Free Tokenization

Abstract
The cost of deploying vision transformers increasingly represents a barrier to wider industrial adoption. Existing compression requires additional end-to-end fine-tuning or incurs a significant drawback to runtime, thus making them ill-suited for online inference. We introduce the Visual Word Tokenizer (VWT), a training-free method for reducing energy costs while retaining performance and runtime. The VWT groups patches (visual subwords) that are frequently used into visual words while infrequent ones remain intact. To do so, intra-image or inter-image statistics are leveraged to identify similar visual concepts for compression. Experimentally, we demonstrate a reduction in wattage of up to 19% with only a 20% increase in runtime at most. Comparative approaches of 8-bit quantization and token merging achieve a lower or similar energy efficiency but exact a higher toll on runtime (up to 2×2$\times$2 × or more). Our results indicate that VWTs are well-suited for efficient online inference with a marginal compromise on performance.
{NoHyper}* Corresponding author: jg717@sussex.ac.uk.
Figure 1: Visualization of patch matching by our visual word tokenizer (${\mathcal{T}}_{inter}^{100}$ model). Patches that are matched with one another are indicated by identical colors. Higher patch matching is exhibited by the background rather than foreground object across the datasets. The patch matching of ${\mathcal{T}}_{inter}^{100}$ serves as a rudimentary form of image segmentation by grouping similar non-adjacent visual concepts.
Figure 2: Overview of the Visual Word Tokenizer. (a) An intra-image approach. In the forward pass, the pixel variance of the patches is computed with bottom-k values being masked. Grouped tokens are dropped after positional information is added. Note that ${\mathcal{V}}=1$ as only a single pixel variance feature is used. (b) An inter-image approach. A Bag-of-Visual Words is formed by clustering patches in the image space. In the forward pass, the minimum pairwise cosine distance between the patches and visual words is computed with values above the threshold being masked. Grouped tokens are averaged after positional information is added.
1 Introduction
In recent years, deep learning has seen continuous integration into a variety of systems worldwide. From coding to gaming, neural networks are increasingly deployed in online scenarios, where asynchronous requests are processed in real-time. However, due to the size and complexity of modern architectures, such models are costly to run in practice. Various methods have been proposed to improve model efficiency such as Knowledge Distillation [22], Pruning [21, 38], and Quantization [11]. However, many of these methods either require end-to-end fine-tuning to recover performance or significantly reduce runtime. In the field of Natural Language Processing (NLP), there is a growing trend towards improving efficiency via tokenization [18, 19, 9, 61, 40]. Newer large language models (LLMs) [53, 14] exhibit a noticeably larger vocabulary than their earlier counterparts [12, 42], thereby producing shorter sequences across various distributions. For computer vision, increasing interest is placed on reducing the cost of deploying the vision transformer (ViT) [13]. As image encoders in larger vision-language systems, ViTs are used to process images as fixed sets of tokens. Similar to downsampling in convolutional neural networks, most research [25, 30, 45, 37, 5, 4, 24] has focused on merging and/or pruning tokens in the intermediate layers to reduce computational overhead. Given the analogous architecture of the transformer across image and text modalities, our work looks instead at the idea of tokenization for efficiency by splitting an image into variable sets of tokens.
To introduce variability, we draw inspiration from subword tokenization algorithms used in NLP which follow the principle that common words should remain intact while infrequent words are broken down into meaningful subword units. Instead of a top-down approach – splitting words into subwords, our work for image data takes a bottom-up approach by grouping visual subwords (image patches) into visual words. We also twist the underlying principle: frequently used image patches should be grouped as they are more likely to describe common features or backgrounds while infrequent ones remain intact as they might carry task-relevant information. We propose two procedures to capture this principle. The first is an intra-image approach where patches with the lowest pixel variance within each image are grouped as they are more likely to represent uniform areas (e.g. backgrounds). The second is an inter-image approach, where basic features across multiple images such as colors or edges are discovered as visual words. Image patches are then grouped based on the similarity of these basic characteristics. Crucially, image patches that have distinct characteristics (i.e. high dissimilarity with any visual word) remain intact and form visual subwords. The code for our paper is publicly available [^1] https://github.com/wearepal/visual-word-tokenizer.
2 Related Work
Efficient ViTs.
Most works for improving the efficiency of ViTs have focused on reducing tokens in the intermediate layers by leveraging importance scores. In [30, 60, 25, 4], redundancy is addressed by fusing tokens. Both [45] and [50] opt to prune such tokens instead. Recent efforts [7, 8, 6, 24] attempt to combine the benefits of merging and pruning. In [54], an additional metric termed the energy score is used to better identify redundancy. Uniquely, [16] use inverse transform sampling to select important tokens. Most relevant to our work are [37] and [5]. The former assigns tokens to centroids via clustering, while the latter progressively merges tokens in a training-free manner.
Specialized Tokenizers.
Our method also takes inspiration from works in NLP for efficient inference. Increasingly, the tokenizer’s vocabulary is specialized to reduce the token length of the given distribution. In [18], domain adaptation of the tokenizer by Fast Vocabulary Transfer (FVT) ensures fewer subword or character tokens are produced. [19] followed up by introducing n-grams for tokenization beyond the word-level boundary. Vocabulary specialization coupled with FVT has also been shown to accelerate code generation for modern LLMs [9]. Meanwhile, [61] analyzed the effectiveness of various vocabulary adaptation techniques for efficient cross-lingual inference. Recently, [40] leveraged hypernetworks for zero-shot tokenizer transfer of newly domain-adapted vocabularies.
Vector Quantization.
The idea of discretizing continuous distributions has been explored in many works, most recently for image generation. [62] leveraged clustering for learning a codebook that maps keypoint descriptors to discrete visual words. In [58, 3], discretization is applied as part of the modeling for ViTs. [55] learned discerete image representations by introducing the Vector Quantised-Variational Autoencoder (VQ-VAE) approach. [15] and [65] further improved upon the VQ-VAE by combining the expressiveness of transformers with an adversarial framework. Increasingly, vision-language models paired with codebooks conduct image synthesis autoregressively [44, 66, 35, 36, 52, 51, 49]. Lastly, [63] tackled disentangled representation learning and scene decomposition by tokenizing images into visual concepts.
3 Visual Word Tokenizer
In ViTs, tokenization is a process that splits an image into patches (tokens) which are then projected into a series of embeddings. The number of patches is typically fixed (e.g. $197$) based on the choice of architecture and hyperparameters. We seek to split the image into variable length inputs instead (i.e. certain images will use $80$ tokens, while others a $100$). This variability is induced by conventional text tokenizers [17, 48, 59, 26] for model efficiency. We propose to achieve this via visual words that group patches (visual subwords) based on some commonality in the image space. These visual words capture frequently used patches while retaining infrequent ones as is. A simple yet effective grouping can be done using either a criterion that looks at statistics of only one image (an intra-image approach) or across many images (an inter-image approach). Figure 2 summarizes the Visual Word Tokenizer (VWT).
3.1 An intra-image approach
The pixel variance of the image patches is the simplest criterion that can be used for grouping. In Figure 2, this approach only utilizes the deploy step by grouping the top-k patches with the lowest pixel variance while leaving the rest separate. To compress the grouped tokens, we opt to drop them as they tend to exhibit excessive homogeneity. We do not include the [CLS] token for dropping. This approach is inspired by [39] which aimed to reduce the training cost of OWLv2, an open-vocabulary object detector. Dropping patches with the lowest pixel variance removes padding areas and uniform backgrounds from the input mosaics of raw images, thus increasing training efficiency.
3.2 An inter-image approach
Inspired by text tokenizers and codebooks for image generation, we propose a variable length tokenizer that statistically discovers visual words across many images. The tokenizer consists of two steps: Pre-Process and Deploy
Pre-Process.
The Bag-of-Visual Words (BoVW) is a popular method for modeling images via discrete representations. In [62], k-means clustering is applied to keypoint descriptors from SIFT [34] to learn a fixed set of centroids. These centroids represent the vocabulary to which multiple descriptors are mapped in a process termed Vector Quantization (VQ). In our method, we adopt a variation of this framework by building the BoVW using patches from the image space. Our design choice is motivated by two main factors. First, we find keypoint descriptors to be costly for inference. In each forward pass, computing keypoints for each image would significantly increase runtime. Second, in our early experimentation, we observed that patches in the embedding space have little similarity to one another. Such issues were also described by [5], thus leading to their use of attention scores instead. Further justification for leveraging the image space is provided in Section B.1.
Given a dataset, we first patchify the images using the same patch size as the image encoder (e.g. 16 for ViT-B/16). We then cluster the patches via k-means to form the BoVW of a given vocabulary size, where each centroid represents a visual word. Patchification is done via basic tensor operations and not the pre-trained convolutional layer of the ViT to avoid projection into embeddings. We also execute this process in batches using the MiniBatchKMeans222from sklearn.cluster import MiniBatchKMeans algorithm due to memory constraints. Note that MiniBatchKMeans uses the Euclidean distance by design. Since clustering is done in the image space, the BoVW may be reused by other ViTs with the same patch size regardless of model type.
Dataset Model $Base$ ${\mathcal{T}}_{intra}^{0.5}$ ${\mathcal{T}}_{inter}^{100}$ ${\mathcal{T}}_{inter}^{1000}$ ${\mathcal{T}}_{inter}^{10000}$
In-Domain ImageNet In-Domain ImageNet In-Domain ImageNet
Waterbirds CLIP 197 99 125 124 144 144 165 169
CelebA 89 88 130 119 163 155
MetaShift 112 109 136 135 162 164
OpenImages (Com.) - 114 - 136 - 162
OpenImages (Rare) - 110 - 133 - 160
COCO BLIP 577 289 267 264 317 312 408 405
NoCaps - 257 - 307 - 403
Table 1: Token length per sample (including [CLS]). ${\mathcal{T}}_{inter}^{{\mathcal{V}}}$ of varying pre-processing data and vocabulary sizes are shown. Unlike text tokenizers, domain specialization (i.e. In-Domain) does not result in greater compression. Smaller vocabularies produce shorter sequences whereas larger vocabularies result in patches being increasingly matched to separate visual words.
Dataset Model $Base$ $Q_{8}$ $ToME$ ${\mathcal{T}}_{intra}^{0.5}$ ${\mathcal{T}}_{inter}^{1000}$
Waterbirds CLIP 123.99 $\pm$ 3.07 73.89 $\pm$ 0.23 91.64 $\pm$ 0.23 110.93 $\pm$ 0.33 107.93 $\pm$ 0.21
CelebA 126.53 $\pm$ 0.75 74.47 $\pm$ 0.18 93.07 $\pm$ 0.13 114.37 $\pm$ 0.85 102.12 $\pm$ 0.20
MetaShift 123.30 $\pm$ 0.60 74.02 $\pm$ 0.43 90.96 $\pm$ 0.33 113.84 $\pm$ 2.40 105.78 $\pm$ 0.15
OpenImages (Com.) 135.05 74.80 94.63 114.62 112.67
OpenImages (Rare) 135.66 75.27 94.04 115.78 110.75
COCO BLIP 191.74 81.68 147.23 169.77 156.25
NoCaps 185.12 81.89 149.26 175.06 155.11
(a) Wattage (watt) $\downarrow$
Dataset Model $Base$ $Q_{8}$ $ToME$ ${\mathcal{T}}_{intra}^{0.5}$ ${\mathcal{T}}_{inter}^{1000}$
Waterbirds CLIP 8.48 $\pm$ 0.11 88.55 $\pm$ 0.10 17.04 $\pm$ 0.06 5.84 $\pm$ 0.03 9.73 $\pm$ 0.03
CelebA 8.33 $\pm$ 0.01 89.41 $\pm$ 3.30 17.25 $\pm$ 0.13 5.86 $\pm$ 0.05 9.54 $\pm$ 0.03
MetaShift 8.21 $\pm$ 0.09 89.48 $\pm$ 0.28 17.23 $\pm$ 0.34 6.01 $\pm$ 0.30 9.87 $\pm$ 0.05
OpenImages (Com.) 8.83 92.23 17.44 5.97 9.69
OpenImages (Rare) 8.60 86.00 17.04 6.19 10.02
COCO BLIP 12.45 57.11 15.06 8.10 9.10
NoCaps 12.44 57.61 15.08 8.04 9.28
(b) Runtime (millisecond) $\downarrow$
Table 2: Wattage and runtime per sample. Compared to $Base$, VWTs reduce wattages by up to 19% with ${\mathcal{T}}_{inter}^{1000}$. Although $Q_{8}$ and $ToME$ display lower wattages than VWTs, they induce significantly longer runtimes, especially the former. ${\mathcal{T}}_{intra}^{0.5}$ is consistently faster than $Base$, while ${\mathcal{T}}_{inter}^{1000}$ may increase runtimes by up to 20% at most (except on COCO and NoCaps).
Deploy.
Once the BoVW is formed, we now turn toward the process of sequence compression. One way of leveraging the BoVW would be to merge similar patches in the image space before projecting them into embeddings. However, such a naive approach will significantly degrade performance as doing so necessitates the initialization of a new linear layer for projection. To avoid this, we begin by patchifying and computing the pairwise cosine distance between the patches and BoVW. For each patch, we retain only the minimum distance. Unlike text, we are unable to obtain exact matches with images. As such, distances higher than a given threshold are masked out to ensure dissimilar patches are not merged. This allows us to ensure that infrequently used image patches remain intact. At this point, we have determined the groupings of similar patches via their connections to the same visual words. We then apply the pre-trained convolutional layer of the ViT on the original image to patchify and project it into a series of embeddings. Before merging, we ensure that positional information is added to the embeddings as we found it to work better than adding them later. Lastly, we average the embeddings element-wise based on the earlier defined groupings. We do not include the [CLS] token for merging.
For the inter-image approach, if batching instead of online inference is desired, the uniform requirement of tensors becomes a challenge. To maintain parallelism, any reduction has to be equal across samples. Due to the non-uniformity of tokenization, sequences have to be sequentially compressed before padding to the same length. We opt to append additional [PAD] tokens until the longest length within the batch is achieved. Similar to text transformers [56], the attention scores are set to negative infinity before the softmax to nullify the influence of padding. We do not add positional information to the [PAD] tokens as extrapolating such information to non-uniform sequences will significantly worsen model efficiency.
4 Experiments
Consider a pre-trained image encoder $f\in{\mathcal{F}}$ with parameters $\theta\in\Theta$ that transforms inputs $x\in{\mathcal{X}}\subseteq\mathbb{R}^{d}$ to encodings $\hat{x}\in\hat{{\mathcal{X}}}\subseteq\mathbb{R}^{d}$. The encodings can then be mapped to labels $y\in{\mathcal{Y}}$ for some given task, be it classification or generation. More specifically, the ViT first transforms inputs $x$ to tokens $t_{1},\ldots,t_{N}$ before further processing by the attention layers. The number of tokens $N$ is a constant defined by $(\frac{I}{P})^{2}$, where $I$ and $P$ are the image and patch sizes, respectively. Let ${\mathcal{T}}$ be the VWT associated with a vocabulary $v\in{\mathcal{V}}$, where $v$ is a visual word learned from some dataset $D$. The tokenizer transforms the input $x$ into tokens $t_{1},\ldots,t_{M}$, where $M\ll N$. In our experiments, we seek to analyze the effect of ${\mathcal{T}}$ on online inference efficiency. We do so through the lens of (i) performance retention in visual recognition tasks of zero-shot classification, (ii) group robustness, and (iii) generative performance of visual captioning.
Model Waterbirds CelebA MetaShift OpenImages
Average $\uparrow$ Worst $\uparrow$ Average $\uparrow$ Worst $\uparrow$ Average $\uparrow$ Worst $\uparrow$ Common $\uparrow$ Rare $\uparrow$
$Base$ 79.06 $\pm$ 0.14 21.86 $\pm$ 0.32 89.61 $\pm$ 0.14 47.21 $\pm$ 2.41 95.31 $\pm$ 1.00 87.69 $\pm$ 3.08 70.48 63.36
$Q_{8}$ 79.94 $\pm$ 0.20 24.25 $\pm$ 0.09 89.73 $\pm$ 0.13 48.19 $\pm$ 3.24 95.23 $\pm$ 0.74 88.21 $\pm$ 3.87 70.48 63.19
$ToME$ 79.80 $\pm$ 0.15 27.05 $\pm$ 0.09 89.25 $\pm$ 0.15 28.52 $\pm$ 4.49 94.24 $\pm$ 0.43 87.37 $\pm$ 3.01 65.69 60.05
${\mathcal{T}}_{intra}^{0.5}$ 75.56 $\pm$ 0.06 31.00 $\pm$ 0.47 89.27 $\pm$ 0.12 50.93 $\pm$ 1.16 92.94 $\pm$ 0.35 86.15 $\pm$ 2.66 65.52 59.73
${\mathcal{T}}_{inter}^{1000}$ 79.19 $\pm$ 0.10 23.83 $\pm$ 0.27 90.79 $\pm$ 0.15 47.96 $\pm$ 4.98 93.94 $\pm$ 0.70 86.67 $\pm$ 2.35 66.15 60.56
(a) CLIP (image classification and group robustness)
Model COCO NoCaps
Karpathy $\uparrow$ In-Domain $\uparrow$ Near-Domain $\uparrow$ Out-of-Domain $\uparrow$ Overall $\uparrow$
BLEU@4 CIDEr CIDEr SPICE CIDEr SPICE CIDEr SPICE CIDEr SPICE
$Base$ 34.00 107.35 103.57 14.60 98.87 14.11 92.96 13.30 98.34 14.02
$Q_{8}$ 33.85 106.87 103.86 14.53 100.07 14.18 92.72 13.49 99.12 14.10
$ToME$ 30.02 94.40 97.35 14.42 90.74 13.39 84.22 12.69 90.37 13.40
${\mathcal{T}}_{intra}^{0.5}$ 32.80 104.37 99.86 14.20 94.21 13.73 89.52 13.12 94.07 13.68
${\mathcal{T}}_{inter}^{1000}$ 32.12 101.79 98.19 13.95 93.85 13.49 86.03 12.67 92.89 13.39
(b) BLIP (image captioning)
Table 3: Zero-shot image classification (CLIP), group robustness (CLIP) and captioning (BLIP). VWTs maintain average accuracy relative to $Base$ while improving robustness on Waterbirds. On COCO and NoCaps, VWTs retains a higher performance than $ToME$.
Dataset and Setting.
To study the effects of tokenization on (i) performance retention in visual recognition tasks of zero-shot classification, we conduct two sets of evaluations: 1) using one dataset for large-scale evaluation and 2) using three datasets for robustness analysis detailed below. We focus on a zero-shot setting by eschewing any end-to-end fine-tuning of $f$. For large-scale evaluation, we use the OpenImages v6 dataset [27]. Following [23], the test split is divided into common and rare subsets that consist of 57 224 and 21 991 images, respectively. The former has 214 classes, while the latter has 200. To perform classification, we compute the cosine similarity between an image embedding and the encoded text label [^3] The prefix ”a photo of a ” is also added to encode each text label..
To study performance retention and (ii) group robustness in image classification, we utilize three publicly available datasets: Waterbirds [57], CelebA [33], and MetaShift [29], that are used as benchmarks in robustness and fairness research [47, 32, 64]. To perform (zero-shot) classification, we compute the cosine similarity between an image embedding and the following encoded text labels3 for Waterbirds, CelebA, and MetaShift, respectively: {’landbird’, ’waterbird’}, {’non-blond’, ’blond’}, {’dog’, ’cat’}. The same data splits and spurious attributes as [32] and [64] are used for the former two and latter datasets, respectively. Specifically, the following spurious attributes are studied: habitation {’water’, ’land’} for Waterbirds, gender {’female’, ’male’} for CelebA, and context {’indoor’, ’outdoor’} for MetaShift. Further details on the subgroups are provided in Section A.1. We also compute the variance by shuffling samples among the data splits while maintaining subgroup proportions. We leverage attribute annotations not only for analyzing average and robust performance but also for the fairness of tokenization in Section B.4.
To analyze the effect of ${\mathcal{T}}$ on (iii) generative task of image captioning, we utilize the Karpathy test split of COCO dataset [31] and a validation set of NoCaps dataset  [1] following the setting in previous work [28].
Finally, to study inference efficiency, we utilize all datasets for the downstream visual tasks (i)-(iii) described above. We focus on the online setting where batch size is 1.
Figure 3: Visualization of the vocabulary of ${\mathcal{T}}_{inter}^{100}$. Patches from ImageNet-1K are matched to the closest visual word using the Euclidean distance. Visual words are shown to depict basic features such as colors or edges. As centroids, each visual word is an average representation of patches that belong to the matched cluster.
Baselines.
For image classification, we load the pre-trained CLIP [43] model from HuggingFace [^4] https://huggingface.co/openai/clip-vit-base-patch16. The prefix ”a photo of a ” is also added to each text category. An image size of 224 $\times$ 224 is used with bilinear interpolation for CLIP. For image captioning, we load the pre-trained BLIP [28] model from HuggingFace [^5] https://huggingface.co/Salesforce/blip-image-captioning-base. To perform zero-shot captioning, we use a beam size of 3 along with maximum and minimum lengths of 20 and 5, respectively. Unlike [28], we do not include the prefix ”a picture of ” as we found it to work better. An image size of 384 $\times$ 384 is used with bicubic interpolation. Both CLIP and BLIP utilize the ViT-B/16 image encoder unless stated otherwise.
Aside from the pre-trained model which we denote as $Base$, we also consider 8-bit quantization [11] and token merging [5] as additional baselines. We denote the former as $Q_{8}$ and the latter as $ToME$. Following [5], we utilize a reduction per layer for $ToME$ of 13 with CLIP and 23 with BLIP due to their respective input image sizes. For the VWTs, we set the top-k of the intra-image approach to 50% of the total number of patches which we denote as ${\mathcal{T}}_{intra}^{0.5}$. For the inter-image approach, we set the threshold to 0.1 unless stated otherwise and denote it as ${\mathcal{T}}_{inter}^{{\mathcal{V}}}$, where ${\mathcal{V}}$ is the size of the vocabulary. Lastly, our experiments are conducted using a single NVIDIA A100 GPU.
Figure 4: Visualization of long-form captioning on COCO. Longer captions are generated via a length penalty of 2.0 and maximum length of 40. Interestingly, the smaller vocabulary of ${\mathcal{T}}_{inter}^{100}$ improves descriptiveness and coherence more than ${\mathcal{T}}_{intra}^{0.5}$ or ${\mathcal{T}}_{inter}^{1000}$.
Inference Efficiency.
First we analyze the effects of VWTs on token length. We seek to understand how the choice of pre-processing data and vocabulary size affects the degree of compression. Table 1 shows the token length per sample (including [CLS]) on different datasets. Unlike ${\mathcal{T}}_{intra}^{0.5}$, the sequence lengths induced by ${\mathcal{T}}_{inter}^{{\mathcal{V}}}$ are not equal. First, we compare ${\mathcal{T}}_{inter}^{{\mathcal{V}}}$ pre-processed on the in-domain dataset and ImageNet-1K [10]. The in-domain dataset is represented by the training split if available. On text data, in-domain tokenizers [9, 61, 40] have been shown to produce shorter sequences by specializing the vocabulary on the given distribution. Interestingly, we observe no such effect with image data as seen by the similar lengths between In-Domain and ImageNet-1K. Only on CelebA, do we see a greater reduction with ${\mathcal{T}}_{inter}^{1000}$ and ${\mathcal{T}}_{inter}^{10000}$ pre-processed on ImageNet-1K. Second, unlike text tokenizers, a trend of decreasing compression is seen as vocabulary size increases. With text, larger vocabularies ensure that more tokens are kept as words rather than subwords. In our case, we posit that an increasing number of patches are matched to separate visual words, thus lowering the overall compression. We restrict our remaining analyses to ${\mathcal{T}}_{inter}^{1000}$ pre-processed on ImageNet-1K unless stated otherwise.
Having analyzed the effects on token length, we turn our attention to the practical metrics of wattage and runtime. For online inference, savings in wattage should not incur a significant cost to runtime. In Table 2, we compare the efficiency of tokenization to $Base$, $Q_{8}$, and $ToME$ on the datasets. First, we find the wattage of VWTs to be lower than $Base$ across the datasets. On CelebA, this reduction is up to 19% with ${\mathcal{T}}_{inter}^{1000}$. Naturally, efficiency decreases as vocabulary size increases due to smaller compression. Although $Q_{8}$ and $ToME$ result in a lower wattage than VWTs, we note the drawbacks in runtime. In particular, $Q_{8}$ results in a significantly longer runtime for $Base$ as tensors need to be repeatedly quantized and dequantized. Meanwhile, $ToME$ can increase runtime by up to 2×2$\times$2 × on Waterbirds, CelebA, and MetaShift. We observe increases of up to 20% at most with ${\mathcal{T}}_{inter}^{1000}$ on the aforementioned datasets. ${\mathcal{T}}_{intra}^{0.5}$ displays lower runtimes than $Base$ across the datasets while ${\mathcal{T}}_{inter}^{1000}$ does so on COCO and NoCaps. We have shown how VWTs are more suitable for online inference by striking a good balance between savings in wattage and costs to runtime.
Performance Retention of Classification, Robustness, and Captioning.
Another important factor in compression is the effect on model performance. In Table 3, we tabulate the performance of image classification and captioning using CLIP [43] and BLIP [28], respectively. For classification and group robustness, we analyze the average and worst-group (robust) performance as done by [47, 46]. First, we find the degradation in average accuracy to be small for ${\mathcal{T}}_{intra}^{0.5}$ and ${\mathcal{T}}_{inter}^{1000}$. The largest drop of up to 2% only is shown by the former on MetaShift. Likewise, $Q_{8}$ and $ToME$ retain a high average accuracy across the datasets. Second, we observe possible improvements in robustness with VWTs. On Waterbirds and CelebA, the worst-group accuracy with ${\mathcal{T}}_{intra}^{0.5}$ increases by up to 29% and 8%, respectively. Only on MetaShift is robustness lower than $Q_{8}$ or $ToME$ relative to $Base$. Like [20], we find compression to not always be harmful to group robustness.
We conduct more evaluations on the large-scale OpenImages v6 dataset for zero-shot classification. We report mean Average Precision (mAP) as done by [23]. From Tables 2 and 3, we find VWTs to be more suitable for online inference than $Q_{8}$ or $ToME$ in the large-scale setting, while retaining comparable performance to the $Base$ model.
For image captioning with BLIP, we evaluate our models following the setting in  [28] by using the BLEU, CIDEr, and SPICE metrics w.r.t. the ground truth captions. On COCO, VWTs display a higher performance than $ToME$ and are competitive with $Q_{8}$. On NoCaps, we see the largest degradation on the out-of-domain samples with ${\mathcal{T}}_{inter}^{1000}$. However, overall performance is still higher than $ToME$. Only $Q_{8}$ displays a slight improvement over $Base$. Like $ToME$, VWTs are shown to not require additional end-to-end fine-tuning for performance retention.
(a) Wattage (watt) $\downarrow$
(b) Runtime (millisecond) $\downarrow$
[…]
Visual Words.
To better understand the inter-image approach, we visualize the patch matching of ${\mathcal{T}}_{inter}^{100}$ in Figure 1. We highlight in identical colors patches that are matched with one another. First, we find that patches representing the background are more frequently grouped than those of the foreground. On Waterbirds, the merging of background patches may explain the improved robustness in Table 3 by mitigating the spurious correlation with the background. On CelebA, ${\mathcal{T}}_{inter}^{100}$ tends to avoid matching the eyes or mouths of the individuals. Second, we observe that patch matching is capable of grouping similar but non-adjacent visual concepts. In certain examples of MetaShift and NoCaps, multiple cats and foods are seemingly matched together, respectively. Our analysis shows that patch matching serves as a rudimentary form of image segmentation.
In Figure 3, we visualize the vocabulary of ${\mathcal{T}}_{inter}^{100}$ to analyze the formation of the visual words. We show patches from ImageNet-1K (i.e. pre-processing data) that are matched to each visual word using the Euclidean distance. Since visual words are centroids, patches that are matched to the same visual word belong to the same cluster. We observe that visual words depict basic features such as colors or edges. These features are also formed as an average representation of the patches that belong to each cluster. By representing basic features, visual words serve as effective anchor points to which patches can be matched to. Analysis on the sparsity of the vocabularies in the inter-image approach can be found in Section B.5.
Long-form Captioning.
The generated captions by BLIP [28] can be used as priors for further training of other models. As such, longer descriptive captions may be more desirable than the typical short captions associated with Internet data. In Figure 4, we visualize the long-form captions on COCO. To enable longer generations, we set the length penalty to 2.0 and double the maximum length to 40. All other generation settings are kept the same. With longer captions, the generation may degenerate into unnecessary repetitions on certain samples. Interestingly, descriptiveness and coherence improves more with ${\mathcal{T}}_{inter}^{100}$ than ${\mathcal{T}}_{intra}^{0.5}$ or ${\mathcal{T}}_{inter}^{1000}$. Additional long-form captions by $Q_{8}$, $ToME$, and ${\mathcal{T}}_{inter}^{10000}$ are provided in Section C.1.
Batch Processing.
Earlier we stated how VWTs works optimally for online inference. In Figure 5, we seek to better understand its effectiveness for offline inference where batch sizes are greater than 1. Using typical batches of $\{4,8,16,32\}$, we compare ${\mathcal{T}}_{intra}^{0.5}$ and ${\mathcal{T}}_{inter}^{1000}$ to $Base$ and $ToME$. First, we find that as batch size increases, lower wattages relative to $Base$ are seen with ${\mathcal{T}}_{intra}^{0.5}$ and $ToME$, particularly at lower batch sizes. Both VWTs and $ToME$ eventually reach wattage parity with $Base$ at larger batches (i.e. 32), with ${\mathcal{T}}_{inter}^{1000}$ doing so much earlier. Second, we observe different runtime trends between VWTs and $ToME$. As batches grow, ${\mathcal{T}}_{inter}^{1000}$ continues to show similar runtimes to $Base$. Meanwhile, ${\mathcal{T}}_{intra}^{0.5}$ and $ToME$ display lower runtimes with the exception of batch size 4 for the latter. As such, during batching, the inter-image approach does not confer the same degree of efficiency improvements.
When to use the intra-image or inter-image approach?
Concerning the choice between the intra-image or inter-image approach, we posit that when global information is required, dropping tokens may be more advantageous by removing unnecessary noise from the visual input. On the other hand, when the task necessitates local information (e.g. long-form captioning, object detection, etc.), merging tokens may better preserve the visual concepts. For example, the image on the first row of Figure 7 shows that ${\mathcal{T}}_{intra}^{0.5}$ removes the mountainous background, thus leading to the absence of ”mountain” in the long-form caption of Figure 4. Potentially, combining both approaches may maximize compression while preserving information by merging the mountains and dropping the bushes in said image.
5 Conclusion
In this work, we set out to define a training-free tokenization for ViTs that lowers wattage while balancing costs to runtime and performance. In online scenarios, we have shown empirically that our intra-image and inter-image approaches are competitive with 8-bit quantization and token merging for image classification and captioning. Qualitatively, we observe how the inter-image approach groups analogous visual concepts via visual words that represent basic features. Analysis on large-scale classification further validates the viability of our method while long-form captioning shows its potential for improving descriptiveness and coherence. Lastly, the inter-image approach has limits for batch processing due to its non-parallelism. As a future work, further research can be made into combining both approaches.
Acknowledgements
This research was supported by a European Research Council (ERC) Starting Grant for the project “Bayesian Models and Algorithms for Fairness and Transparency”, funded under the European Union’s Horizon 2020 Framework Programme (grant agreement no. 851538). Novi Quadrianto is also supported by the Basque Government through the BERC 2022-2025 program and by the Ministry of Science and Innovation: BCAM Severo Ochoa accreditation CEX2021-001142-S / MICIN/ AEI/ 10.13039/501100011033. Viktoriia Sharmanska is currently at Epic Games.
Appendix A Further Details
A.1 Datasets for Group Robustness
Here, we detail the task of each dataset for group robustness. In Table 4, we also tabulate the labels and attributes that define each subgroup along with their sample sizes.
Waterbirds.
Given an image of a bird, the task is to predict whether it is a waterbird or landbird [57]. The attribute is the background that the bird is on [47].
CelebA.
Given an image of a person, the task is to predict whether their hair color is blond or not [33]. The attribute is the binary gender of the person.
MetaShift.
Given an image of an animal, the task is to predict whether it is a dog or cat [29]. The attribute is the environment that the dog or cat is in.
Dataset Subgroup Label Attribute Training Validation Test
Waterbirds 0 0 (landbird) 0 (on land) 3498 467 2255
1 0 (landbird) 1 (on water) 184 466 2255
2 1 (waterbird) 0 (on land) 56 133 642
3 1 (waterbird) 1 (on water) 1057 133 642
CelebA 0 0 (non-blond) 0 (woman) 71629 8535 9767
1 0 (non-blond) 1 (man) 66874 8276 7535
2 1 (blond) 0 (woman) 22880 2874 2480
3 1 (blond) 1 (man) 1387 182 180
MetaShift 0 0 (dog) 0 (outdoor) 784 127 273
1 0 (dog) 1 (indoor) 507 75 191
2 1 (cat) 0 (outdoor) 196 33 65
3 1 (cat) 1 (indoor) 789 114 345
Table 4: Defined subgroups in Waterbirds, CelebA, and MetaShift.
Appendix B Supplementary Experiments
Here, we detail the supplementary experiments that we conducted. First, we explore additional design choices and effects of the inter-image approach. Section B.1 shows the ineffectiveness of patch matching with visual words formed in the embedding space. In Section B.2, random matching is applied to further determine if matching by the visual words are meaningful. Section B.3 ablates the similarity threshold by relaxing it for further compression. The fairness of tokenization and sparsity of the vocabulary are shown in Sections B.4 and B.5, respectively.
B.1 Ineffectiveness of the Embedding Space
In Table 5, we analyze the sequence compression using visual words formed in the embedding space. To do so, we initialize each vocabulary by clustering embeddings from the pre-trained convolution layer of CLIP or BLIP. During inference, we match the activations from the pre-trained convolution layer with the visual words. Compared to Table 1, we observe a notable reduction in the degree of compression irrespective of pre-processing data and vocabulary size. In [50], embeddings are shown to become progressively more similar to one another due to self-attention. Hence, it is unsurprising that matching activations and visual words at the beginning of the network is ineffective.
B.2 Random Matching of Visual Words
In Table 6, we study the effects of randomly matching the patches. We initialize the pairwise cosine distance by sampling from a uniform distribution of $[0,2]$. First, we find the token length (including [CLS]) to differ noticeably from Table 1. For ${\mathcal{T}}_{inter}^{100}$, sequences are further reduced across the datasets. Conversely, ${\mathcal{T}}_{inter}^{1000}$ and ${\mathcal{T}}_{inter}^{10000}$ display little compression. Second, performance is shown to change from Table 3. We observe a significant degradation in average and worst-group accuracies with ${\mathcal{T}}_{inter}^{100}$ on Waterbirds and MetaShift. For ${\mathcal{T}}_{inter}^{1000}$ and ${\mathcal{T}}_{inter}^{10000}$, performance does not shift much from $Base$ as barely any compression occurs. With random matching, the captions become completely different than those of $Base$ in Figure 10.
B.3 Ablation of the Similarity Threshold
We have shown how VWTs can improve the efficiency of online inference. To better understand their limitations, we ablate the threshold of ${\mathcal{T}}_{inter}^{100}$ by setting it to values of $\{0.2,0.3,0.4,0.5\}$ in Table 7. We seek to determine if exploiting higher thresholds for increased compression is a viable strategy. Naturally, we observe a reduction in performance as increasingly dissimilar patches are merged. For Waterbirds and MetaShift, the worst-group accuracy degrades more significantly than the average, especially with the former. Interestingly, average accuracy remains relatively unchanged while worst-group accuracy improves significantly on CelebA irrespective of similarity threshold. We posit that at higher thresholds, the merging of core features represented by the foreground object results in the reduced performance of Waterbirds and MetaShift.
B.4 Fairness of the Tokenization
On text data, tokenizers have been shown to induce unfairness between languages [41, 2]. Since most LLMs are built using English data, tokenization of minority languages produces inherently longer sequences that raises compute costs. We seek to analyze if similar effects exist with VWTs as well. In Table 8, we show the breakdown in token length (including [CLS]) and accuracy (w.r.t $Base$) by subgroup. First, we observe a notable difference in compression between the subgroups of Waterbirds. With ${\mathcal{T}}_{100}$, sequences might differ by up to 39 tokens as seen with subgroups 0 and 3. Smaller discrepancies are displayed on CelebA and MetaShift except for ${\mathcal{T}}_{100}$ on the former. Second, we find that compression does not affect all subgroups equally. Accuracy improves on certain subgroups and degrades on others. A stronger compression does not correlate also with a large change in performance. Like text tokenizers, VWTs may induce inequality in compression that leads to varying inference costs and performance for different subgroups.
B.5 Sparsity of the Vocabulary
To better understand the utilization of the visual words, we plot the probability distribution of the matches in Figure 6. Regardless of the dataset, we find that certain visual words are matched more frequently than others, thus leading to a large skew in the distributions. Greater sparsity is also displayed by larger vocabularies as many visual words remain unused across datasets. As such, pruning may be applied to achieve a more efficient vocabulary size post-clustering.
Appendix C Additional Results
Here, we provide additional results by other models and hyperparameters in our experiments. Section C.1 lists long-form captions by $Q_{8}$, $ToME$, and ${\mathcal{T}}_{inter}^{10000}$. Sections C.2 and C.3 detail the intra-image and inter-image approaches with various dropping ratios and vocabulary sizes, respectively.
C.1 Long-form Captions
In Figure 7, we showcase long-form captions by $Q_{8}$, $ToME$, and ${\mathcal{T}}_{inter}^{10000}$. First, we observe improvements in the accuracy of the descriptions with $Q_{8}$ over $Base$ in Figure 4. Although the captions relate better to the image, the overall length is shorter than those of other methods. The generation may continue to introduce unnecessary repetitions. Second, we find $ToME$ and ${\mathcal{T}}_{inter}^{10000}$ to also be beneficial for descriptiveness and coherence. On certain samples, repeating n-grams may still occur.
C.2 Dropping Ratio
Following [39], we apply the intra-image approach with ratios of $\{0.25,0.33,0.7\}$ to the total number of patches in Table 9. We observe a natural degradation as the ratio increases with performance dropping steeply at 0.7. In Figures 8 and 9, we visualize the dropped patches and captions on COCO, respectively. In most cases, the patches with the lowest pixel variance correspond to uninformative backgrounds. Dropping such patches may continue to preserve the foreground object.
C.3 Vocabulary Size
In Table 10, we apply the inter-image approach with vocabulary sizes of $\{100,10000\}$. We find performance to be noticeably lower than $Base$ in Table 3 with the smaller vocabulary of ${\mathcal{T}}_{inter}^{100}$. Meanwhile, performance retention is higher for ${\mathcal{T}}_{inter}^{10000}$ due to lower sequence compression in Table 1. When visualizing the captions on COCO in Figure 10, we also observe that increasing thresholds cause larger deviations from $Base$ than reducing the vocabulary size.
Dataset ${\mathcal{T}}_{inter}^{100}$ ${\mathcal{T}}_{inter}^{1000}$ ${\mathcal{T}}_{inter}^{10000}$
In-Domain ImageNet In-Domain ImageNet In-Domain ImageNet
Waterbirds 180 179 179 179 182 182
CelebA 153 154 157 154 170 165
MetaShift 176 173 175 174 180 180
COCO 405 406 409 408 442 440
NoCaps - 394 - 396 - 431
Table 5: Token length per sample (including [CLS]). Tokenizers of varying pre-processing data and vocabulary sizes are shown. Visual words are formed from embeddings in the pre-trained convolution layer of CLIP or BLIP and not the image patches. Poor compression is seen irrespective of pre-processing data and vocabulary size due to the lack of similarity between the activations and visual words.
Dataset Model ${\mathcal{T}}_{intra}^{100}$ ${\mathcal{T}}_{intra}^{1000}$ ${\mathcal{T}}_{intra}^{10000}$
Waterbirds CLIP 88 179 195
CelebA
MetaShift
OpenImages (Com.)
OpenImages (Rare)
COCO BLIP 104 439 561
NoCaps
(a) Length
Model Waterbirds CelebA MetaShift Overall $\uparrow$
Average $\uparrow$ Worst $\uparrow$ Average $\uparrow$ Worst $\uparrow$ Average $\uparrow$ Worst $\uparrow$ Common $\uparrow$ Rare $\uparrow$
${\mathcal{T}}_{inter}^{100}$ 66.60 $\pm$ 0.28 9.61 $\pm$ 1.51 90.24 $\pm$ 0.06 56.48 $\pm$ 4.32 76.16 $\pm$ 0.70 61.43 $\pm$ 1.32 41.06 39.84
${\mathcal{T}}_{inter}^{1000}$ 78.10 $\pm$ 0.18 22.74 $\pm$ 1.24 90.25 $\pm$ 0.15 47.96 $\pm$ 2.24 94.20 $\pm$ 0.63 87.18 $\pm$ 1.78 70.29 63.48
${\mathcal{T}}_{inter}^{10000}$ 78.98 $\pm$ 0.07 21.96 $\pm$ 0.31 89.61 $\pm$ 0.07 47.46 $\pm$ 2.62 95.27 $\pm$ 0.97 87.18 $\pm$ 2.35 70.52 63.40
(b) CLIP (image classification and group robustness)
Model COCO NoCaps
Karpathy $\uparrow$ In-Domain $\uparrow$ Near-Domain $\uparrow$ Out-of-Domain $\uparrow$ Overall $\uparrow$
BLEU@4 CIDEr CIDEr SPICE CIDEr SPICE CIDEr SPICE CIDEr SPICE
${\mathcal{T}}_{inter}^{100}$ 9.92 19.18 15.81 7.20 10.95 6.25 10.55 6.63 11.57 6.47
${\mathcal{T}}_{inter}^{1000}$ 32.06 100.52 96.87 14.04 91.83 13.47 89.52 12.96 92.09 13.45
${\mathcal{T}}_{inter}^{10000}$ 34.04 107.35 103.55 14.49 98.22 14.00 94.08 13.48 98.15 13.97
(c) BLIP (image captioning)
Table 6: Token length (including [CLS]) and performance of VWTs with random matching. The pairwise cosine distance is initialized by sampling from a uniform distribution of $[0,2]$. Average and worst-group accuracies degrade noticeably with ${\mathcal{T}}_{100}$ except for CelebA. Performance does not change much for ${\mathcal{T}}_{1000}$ and ${\mathcal{T}}_{10000}$ from $Base$ since barely any compression occurs.
Dataset Thresh Length Average $\uparrow$ Worst $\uparrow$
Waterbirds 0.2 97 76.71 $\pm$ 0.24 24.09 $\pm$ 0.59
0.3 79 74.85 $\pm$ 0.15 21.13 $\pm$ 0.36
0.4 66 73.35 $\pm$ 0.12 16.30 $\pm$ 0.24
0.5 58 72.81 $\pm$ 0.10 13.14 $\pm$ 0.48
CelebA 0.2 66 89.44 $\pm$ 0.08 56.48 $\pm$ 1.70
0.3 55 88.85 $\pm$ 0.26 58.70 $\pm$ 4.21
0.4 50 88.31 $\pm$ 0.14 57.04 $\pm$ 3.26
0.5 48 88.21 $\pm$ 0.32 55.74 $\pm$ 1.40
MetaShift 0.2 84 86.16 $\pm$ 0.89 77.46 $\pm$ 2.34
0.3 69 81.16 $\pm$ 1.00 67.77 $\pm$ 1.43
0.4 60 77.73 $\pm$ 1.03 60.65 $\pm$ 2.57
0.5 55 74.75 $\pm$ 1.36 56.08 $\pm$ 3.38
Table 7: Token length (including [CLS]) and performance using ${\mathcal{T}}_{inter}^{100}$ with varying similarity thresholds. Higher thresholds can reduce sequences to 48 tokens on CelebA. Greater compression results in higher degradation in performance, particularly in robustness.
Model Subgroup Waterbirds CelebA MetaShift
Length $\Delta$ Accuracy Length $\Delta$ Accuracy Length $\Delta$ Accuracy
$Base$ 0 197 98.89 197 95.63 197 98.54
1 82.45 96.23 93.19
2 21.86 48.67 87.69
3 54.73 50.00 95.36
${\mathcal{T}}_{inter}^{100}$ 0 144 -0.67 85 -3.25 118 -0.74
1 106 -3.82 88 -0.13 110 -6.18
2 140 19.00 98 32.62 119 -2.92
3 105 6.17 99 6.30 100 -9.02
${\mathcal{T}}_{inter}^{1000}$ 0 160 -0.40 119 -0.46 138 -0.25
1 129 -0.22 117 0.74 138 -2.06
2 160 9.03 124 18.92 143 -1.17
3 129 2.18 124 -4.07 131 -2.13
${\mathcal{T}}_{inter}^{10000}$ 0 180 -0.09 158 0.25 163 -0.37
1 156 1.34 151 0.66 167 0.19
2 181 4.75 158 2.96 169 -1.75
3 158 1.71 156 -6.30 163 -0.81
Table 8: Distribution of token length (including [CLS]) and accuracy (w.r.t. $Base$) by subgroup. Tokenizers of varying vocabulary sizes are shown. Like text tokenizers, VWTs may induce significantly unequal token lengths as seen with ${\mathcal{T}}_{inter}^{100}$ on Waterbirds. Subgroups are also affected unequally as stronger compression does not correlate with greater improvement or degradation in performance.
(a) Waterbirds
(b) CelebA
(c) MetaShift
Figure 6: Probability distribution of the matched visual words. Tokenizers of varying vocabulary sizes are shown. The probability distribution exhibits a large skew irrespective of the dataset as certain visual words are matched more frequently than others. Larger vocabularies display greater sparsity as the many visual words that remain unmatched may be pruned.
Figure 7: Visualization of long-form captions by $Q_{8}$, $ToME$, and ${\mathcal{T}}_{inter}^{10000}$ on COCO.
Dataset Model ${\mathcal{T}}_{intra}^{0.25}$ ${\mathcal{T}}_{intra}^{0.33}$ ${\mathcal{T}}_{intra}^{0.7}$
Waterbirds CLIP 148 132 59
CelebA
MetaShift
OpenImages (Com.)
OpenImages (Rare)
COCO BLIP 433 386 173
NoCaps
(a) Length
Model Waterbirds CelebA MetaShift OpenImages
Average $\uparrow$ Worst $\uparrow$ Average $\uparrow$ Worst $\uparrow$ Average $\uparrow$ Worst $\uparrow$ Common $\uparrow$ Rare $\uparrow$
${\mathcal{T}}_{intra}^{0.25}$ 78.31 $\pm$ 0.09 26.69 $\pm$ 0.24 89.46 $\pm$ 0.16 48.47 $\pm$ 1.56 94.70 $\pm$ 0.43 87.69 $\pm$ 1.54 69.79 63.37
${\mathcal{T}}_{intra}^{0.33}$ 77.72 $\pm$ 0.24 27.78 $\pm$ 0.48 89.31 $\pm$ 0.22 48.83 $\pm$ 1.42 94.62 $\pm$ 0.41 87.18 $\pm$ 2.35 69.11 62.75
${\mathcal{T}}_{intra}^{0.7}$ 73.27 $\pm$ 0.12 18.80 $\pm$ 0.32 89.19 $\pm$ 0.05 45.37 $\pm$ 1.28 82.61 $\pm$ 1.10 63.59 $\pm$ 0.89 46.83 42.68
(b) CLIP (image classification and group robustness)
Model COCO NoCaps
Karpathy $\uparrow$ In-Domain $\uparrow$ Near-Domain $\uparrow$ Out-of-Domain $\uparrow$ Overall $\uparrow$
BLEU@4 CIDEr CIDEr SPICE CIDEr SPICE CIDEr SPICE CIDEr SPICE
${\mathcal{T}}_{intra}^{0.25}$ 33.92 107.47 103.59 14.65 98.39 14.07 94.47 13.59 98.34 14.06
${\mathcal{T}}_{intra}^{0.33}$ 33.59 106.25 103.19 14.65 97.40 13.95 92.66 13.41 97.27 13.95
${\mathcal{T}}_{intra}^{0.7}$ 29.64 93.20 89.89 13.53 85.04 12.82 80.81 12.31 84.88 12.82
(c) BLIP (image captioning)
Table 9: Token length (including [CLS]) and performance of VWTs with varying dropping ratios.
Figure 8: Visualization of patch dropping by our visual word tokenizer (${\mathcal{T}}_{intra}^{0.5}$ model). Patches with the lowest pixel variance that are dropped are indicated in black. In most cases, the dropped patches correspond to the background object that is typically uninformative.
Figure 9: Visualization of captions by token dropping on COCO. Token dropping of up to 0.5 may not cause significant deviation in the generations from the base model. On certain samples, the accuracy of the descriptions can improve as well.
Model Waterbirds CelebA MetaShift OpenImages
Average $\uparrow$ Worst $\uparrow$ Average $\uparrow$ Worst $\uparrow$ Average $\uparrow$ Worst $\uparrow$ Common $\uparrow$ Rare $\uparrow$
${\mathcal{T}}_{inter}^{100}$ 78.41 $\pm$ 0.09 26.01 $\pm$ 0.41 90.04 $\pm$ 0.22 53.15 $\pm$ 3.21 90.24 $\pm$ 0.40 84.28 $\pm$ 2.38 62.90 58.10
${\mathcal{T}}_{inter}^{10000}$ 79.68 $\pm$ 0.11 22.90 $\pm$ 0.16 90.12 $\pm$ 0.18 46.85 $\pm$ 4.63 94.81 $\pm$ 0.95 86.15 $\pm$ 1.54 69.03 62.50
(a) CLIP (image classification and group robustness)
Model COCO NoCaps
Karpathy $\uparrow$ In-Domain $\uparrow$ Near-Domain $\uparrow$ Out-of-Domain $\uparrow$ Overall $\uparrow$
BLEU@4 CIDEr CIDEr SPICE CIDEr SPICE CIDEr SPICE CIDEr SPICE
${\mathcal{T}}_{inter}^{100}$ 31.10 97.70 95.58 13.78 89.62 13.11 82.42 12.44 89.02 13.08
${\mathcal{T}}_{inter}^{10000}$ 33.18 105.08 102.84 14.59 98.03 13.95 91.52 13.14 97.40 13.88
(b) BLIP (image captioning)
Table 10: Token length (including [CLS]) and performance of VWTs with varying vocabulary sizes.
Figure 10: Visualization of captions by VWTs on COCO. Increasing the similarity threshold causes the generated captions to deviate more than reducing the vocabulary size. With random matching, the model begins to completely misunderstand the image.

### v2

@@
 The cost of deploying vision transformers increasingly represents a barrier to wider industrial adoption.
-Existing compression requires additional end-to-end fine-tuning or incurs a significant drawback to runtime, thus making them ill-suited for online inference.
+Existing compression techniques require additional end-to-end fine-tuning or incur a significant drawback to runtime, making them ill-suited for online (real-time) inference, where a prediction is made on any new input as it comes in.
 We introduce the Visual Word Tokenizer (VWT), a training-free method for reducing energy costs while retaining performance and runtime.
-The VWT groups patches (visual subwords) that are frequently used into visual words while infrequent ones remain intact.
-To do so, intra-image or inter-image statistics are leveraged to identify similar visual concepts for compression.
-Experimentally, we demonstrate a reduction in wattage of up to 19% with only a 20% increase in runtime at most.
-Comparative approaches of 8-bit quantization and token merging achieve a lower or similar energy efficiency but exact a higher toll on runtime (up to 2×2$\times$2 × or more).
+The VWT groups visual subwords (image patches) that are frequently used into visual words while infrequent ones remain intact.
+To do so, intra-image or inter-image statistics are leveraged to identify similar visual concepts for sequence compression.
+Experimentally, we demonstrate a reduction in wattage of up to 25% with only a 20% increase in runtime at most.
+Comparative approaches of 8-bit quantization and token merging achieve a lower or similar energy efficiency but exact a higher toll on runtime (up to 100% or more).
 Our results indicate that VWTs are well-suited for efficient online inference with a marginal compromise on performance.
@@
 Patches that are matched with one another are indicated by identical colors.
-Higher patch matching is exhibited by the background rather than foreground object across the datasets.
-The patch matching of ${\mathcal{T}}_{inter}^{100}$ serves as a rudimentary form of image segmentation by grouping similar non-adjacent visual concepts.
-Figure 2: Overview of the Visual Word Tokenizer.
-(a) An intra-image approach.
-In the forward pass, the pixel variance of the patches is computed with bottom-k values being masked.
-Grouped tokens are dropped after positional information is added.
-Note that ${\mathcal{V}}=1$ as only a single pixel variance feature is used.
-(b) An inter-image approach.
-A Bag-of-Visual Words is formed by clustering patches in the image space.
-In the forward pass, the minimum pairwise cosine distance between the patches and visual words is computed with values above the threshold being masked.
-Grouped tokens are averaged after positional information is added.
+Higher patch matching is exhibited by the background rather than foreground across datasets.
+Patch matching serves as a rudimentary form of image segmentation by grouping similar non-adjacent visual concepts.
 1 Introduction
 In recent years, deep learning has seen continuous integration into a variety of systems worldwide.
-From coding to gaming, neural networks are increasingly deployed in online scenarios, where asynchronous requests are processed in real-time.
+From coding to gaming, neural networks are increasingly deployed in online scenarios where asynchronous requests are processed in real-time.
 However, due to the size and complexity of modern architectures, such models are costly to run in practice.
-Various methods have been proposed to improve model efficiency such as Knowledge Distillation [22], Pruning [21, 38], and Quantization [11].
+To address this, various methods have been proposed to improve model efficiency such as Knowledge Distillation [24], Pruning [22, 41], and Quantization [11].
 However, many of these methods either require end-to-end fine-tuning to recover performance or significantly reduce runtime.
@@
 Given the analogous architecture of the transformer across image and text modalities, our work looks instead at the idea of tokenization for efficiency by splitting an image into variable sets of tokens.
-To introduce variability, we draw inspiration from subword tokenization algorithms used in NLP which follow the principle that common words should remain intact while infrequent words are broken down into meaningful subword units.
+To introduce variability, we draw inspiration from subword tokenization algorithms [17, 51] used in NLP which follow the principle that common words should remain intact while infrequent ones are broken down into meaningful subword units.
 Instead of a top-down approach – splitting words into subwords, our work for image data takes a bottom-up approach by grouping visual subwords (image patches) into visual words.
-We also twist the underlying principle: frequently used image patches should be grouped as they are more likely to describe common features or backgrounds while infrequent ones remain intact as they might carry task-relevant information.
+We also twist the underlying principle: frequently used patches should be grouped as they are more likely to describe common features while infrequent ones remain intact as they might carry task-relevant information.
 We propose two procedures to capture this principle.
-The first is an intra-image approach where patches with the lowest pixel variance within each image are grouped as they are more likely to represent uniform areas (e.g.
+The first is an intra-image approach where patches with the lowest pixel variance within each image are grouped as they typically represent uniform areas (e.g.
 backgrounds).
-The second is an inter-image approach, where basic features across multiple images such as colors or edges are discovered as visual words.
+The second is an inter-image approach where basic features across multiple images such as colors or edges are discovered as visual words.
 Image patches are then grouped based on the similarity of these basic characteristics.
-Crucially, image patches that have distinct characteristics (i.e.
-high dissimilarity with any visual word) remain intact and form visual subwords.
+Crucially, patches that have distinct characteristics (i.e.
+high dissimilarity with any visual word) remain intact and form separate visual subwords.
 The code for our paper is publicly available [^1] https://github.com/wearepal/visual-word-tokenizer.
@@
 Most works for improving the efficiency of ViTs have focused on reducing tokens in the intermediate layers by leveraging importance scores.
-In [30, 60, 25, 4], redundancy is addressed by fusing tokens.
-Both [45] and [50] opt to prune such tokens instead.
+In Liang et al.
+[32], Xu et al.
+[66], Kong et al.
+[27], Bian et al.
+[4], redundancy is addressed by fusing tokens.
+Both Rao et al.
+[48] and Tang et al.
+[55] opt to prune such tokens instead.
 Recent efforts [7, 8, 6, 26] attempt to combine the benefits of merging and pruning.
-In [54], an additional metric termed the energy score is used to better identify redundancy.
-Uniquely, [16] use inverse transform sampling to select important tokens.
-Most relevant to our work are [37] and [5].
-The former assigns tokens to centroids via clustering, while the latter progressively merges tokens in a training-free manner.
+In Tran et al.
+[60], an additional metric termed the energy score is used to better identify redundancy.
+Uniquely, Fayyaz et al.
+[16] use inverse transform sampling to select important tokens.
+Most relevant to our work are Marin et al.
+[40] and Bolya et al.
+[5].
+The former assigns tokens to centroids via clustering, while the latter progressively merges tokens layer-by-layer in a training-free manner [^2] Unlike Bolya et al.
+[5], we do not include Marin et al.
+[40] as one of our baselines due to a lack of code release..
+Figure 2: Overview of the Visual Word Tokenizer.
+An intra-image approach.
+In the forward pass, the pixel variance of the patches is computed with the top-k lowest values being masked (deploy).
+The grouped tokens are dropped after positional information is added.
+Note that ${\mathcal{V}}=1$ as only a single pixel variance feature is used.
+An inter-image approach.
+A Bag-of-Visual Words is formed by clustering patches in the pixel space (pre-process).
+In the forward pass, the minimum pairwise cosine distance between the patches and visual words is computed with values above the threshold being masked (deploy).
+The grouped tokens are averaged after positional information is added.
 Specialized Tokenizers.
-Our method also takes inspiration from works in NLP for efficient inference.
-Increasingly, the tokenizer’s vocabulary is specialized to reduce the token length of the given distribution.
-In [18], domain adaptation of the tokenizer by Fast Vocabulary Transfer (FVT) ensures fewer subword or character tokens are produced.
+Our method also takes inspiration from efficient inference in NLP.
+Increasingly, the tokenizer’s vocabulary is specialized to reduce the input token length.
+In Gee et al.
+[18], domain adaptation of the tokenizer ensures fewer subword or character tokens are produced.
+Gee et al.
 [19] followed up by introducing n-grams for tokenization beyond the word-level boundary.
-Vocabulary specialization coupled with FVT has also been shown to accelerate code generation for modern LLMs [9].
-Meanwhile, [61] analyzed the effectiveness of various vocabulary adaptation techniques for efficient cross-lingual inference.
-Recently, [40] leveraged hypernetworks for zero-shot tokenizer transfer of newly domain-adapted vocabularies.
-Vector Quantization.
-The idea of discretizing continuous distributions has been explored in many works, most recently for image generation.
-[62] leveraged clustering for learning a codebook that maps keypoint descriptors to discrete visual words.
-In [58, 3], discretization is applied as part of the modeling for ViTs.
-[55] learned discerete image representations by introducing the Vector Quantised-Variational Autoencoder (VQ-VAE) approach.
-[15] and [65] further improved upon the VQ-VAE by combining the expressiveness of transformers with an adversarial framework.
-Increasingly, vision-language models paired with codebooks conduct image synthesis autoregressively [44, 66, 35, 36, 52, 51, 49].
-Lastly, [63] tackled disentangled representation learning and scene decomposition by tokenizing images into visual concepts.
-3 Visual Word Tokenizer
-In ViTs, tokenization is a process that splits an image into patches (tokens) which are then projected into a series of embeddings.
-The number of patches is typically fixed (e.g.
-$197$) based on the choice of architecture and hyperparameters.
-We seek to split the image into variable length inputs instead (i.e.
-certain images will use $80$ tokens, while others a $100$).
-This variability is induced by conventional text tokenizers [17, 48, 59, 26] for model efficiency.
-We propose to achieve this via visual words that group patches (visual subwords) based on some commonality in the image space.
-These visual words capture frequently used patches while retaining infrequent ones as is.
-A simple yet effective grouping can be done using either a criterion that looks at statistics of only one image (an intra-image approach) or across many images (an inter-image approach).
-Figure 2 summarizes the Visual Word Tokenizer (VWT).
-3.1 An intra-image approach
-The pixel variance of the image patches is the simplest criterion that can be used for grouping.
-In Figure 2, this approach only utilizes the deploy step by grouping the top-k patches with the lowest pixel variance while leaving the rest separate.
-To compress the grouped tokens, we opt to drop them as they tend to exhibit excessive homogeneity.
-We do not include the [CLS] token for dropping.
-This approach is inspired by [39] which aimed to reduce the training cost of OWLv2, an open-vocabulary object detector.
-Dropping patches with the lowest pixel variance removes padding areas and uniform backgrounds from the input mosaics of raw images, thus increasing training efficiency.
-3.2 An inter-image approach
-Inspired by text tokenizers and codebooks for image generation, we propose a variable length tokenizer that statistically discovers visual words across many images.
-The tokenizer consists of two steps: Pre-Process and Deploy
-Pre-Process.
-The Bag-of-Visual Words (BoVW) is a popular method for modeling images via discrete representations.
-In [62], k-means clustering is applied to keypoint descriptors from SIFT [34] to learn a fixed set of centroids.
-These centroids represent the vocabulary to which multiple descriptors are mapped in a process termed Vector Quantization (VQ).
-In our method, we adopt a variation of this framework by building the BoVW using patches from the image space.
-Our design choice is motivated by two main factors.
-First, we find keypoint descriptors to be costly for inference.
-In each forward pass, computing keypoints for each image would significantly increase runtime.
-Second, in our early experimentation, we observed that patches in the embedding space have little similarity to one another.
-Such issues were also described by [5], thus leading to their use of attention scores instead.
-Further justification for leveraging the image space is provided in Section B.1.
-Given a dataset, we first patchify the images using the same patch size as the image encoder (e.g.
-16 for ViT-B/16).
-We then cluster the patches via k-means to form the BoVW of a given vocabulary size, where each centroid represents a visual word.
-Patchification is done via basic tensor operations and not the pre-trained convolutional layer of the ViT to avoid projection into embeddings.
-We also execute this process in batches using the MiniBatchKMeans222from sklearn.cluster import MiniBatchKMeans algorithm due to memory constraints.
-Note that MiniBatchKMeans uses the Euclidean distance by design.
-Since clustering is done in the image space, the BoVW may be reused by other ViTs with the same patch size regardless of model type.
+In Dagan et al.
+[9], tokenizer specialization is also shown to accelerate the task of code generation with modern LLMs.
+Meanwhile, Yamaguchi et al.
+[67] analyzed the effectiveness of various vocabulary adaptation techniques for efficient cross-lingual inference.
+Recently, Minixhofer et al.
+[43] leveraged hypernetworks for zero-shot tokenizer transfer of newly domain-adapted vocabularies.
 Dataset Model $Base$ ${\mathcal{T}}_{intra}^{0.5}$ ${\mathcal{T}}_{inter}^{100}$ ${\mathcal{T}}_{inter}^{1000}$ ${\mathcal{T}}_{inter}^{10000}$
@@
 Smaller vocabularies produce shorter sequences whereas larger vocabularies result in patches being increasingly matched to separate visual words.
-Dataset Model $Base$ $Q_{8}$ $ToME$ ${\mathcal{T}}_{intra}^{0.5}$ ${\mathcal{T}}_{inter}^{1000}$
-Waterbirds CLIP 123.99 $\pm$ 3.07 73.89 $\pm$ 0.23 91.64 $\pm$ 0.23 110.93 $\pm$ 0.33 107.93 $\pm$ 0.21
-CelebA 126.53 $\pm$ 0.75 74.47 $\pm$ 0.18 93.07 $\pm$ 0.13 114.37 $\pm$ 0.85 102.12 $\pm$ 0.20
-MetaShift 123.30 $\pm$ 0.60 74.02 $\pm$ 0.43 90.96 $\pm$ 0.33 113.84 $\pm$ 2.40 105.78 $\pm$ 0.15
-OpenImages (Com.) 135.05 74.80 94.63 114.62 112.67
-OpenImages (Rare) 135.66 75.27 94.04 115.78 110.75
-COCO BLIP 191.74 81.68 147.23 169.77 156.25
-NoCaps 185.12 81.89 149.26 175.06 155.11
+Dataset Model $Base$ $Q_{8}$ $ToME$ ${\mathcal{T}}_{intra}^{0.5}$ ${\mathcal{T}}_{inter}^{100}$ ${\mathcal{T}}_{inter}^{1000}$ ${\mathcal{T}}_{inter}^{10000}$
+Waterbirds CLIP 123.99 73.89 91.64 110.93 102.65 107.93 117.36
+CelebA 126.53 74.47 93.07 114.37 96.31 102.12 115.46
+MetaShift 123.30 74.02 90.96 113.84 98.98 105.78 114.70
+OpenImages (Com.) 135.05 74.80 94.63 114.62 106.33 112.67 123.95
+OpenImages (Rare) 135.66 75.27 94.04 115.78 104.73 110.75 123.05
+COCO BLIP 191.74 81.68 147.23 169.77 144.71 156.25 171.84
+NoCaps 185.12 81.89 149.26 175.06 141.20 155.11 170.38
 (a) Wattage (watt) $\downarrow$
-Dataset Model $Base$ $Q_{8}$ $ToME$ ${\mathcal{T}}_{intra}^{0.5}$ ${\mathcal{T}}_{inter}^{1000}$
-Waterbirds CLIP 8.48 $\pm$ 0.11 88.55 $\pm$ 0.10 17.04 $\pm$ 0.06 5.84 $\pm$ 0.03 9.73 $\pm$ 0.03
-CelebA 8.33 $\pm$ 0.01 89.41 $\pm$ 3.30 17.25 $\pm$ 0.13 5.86 $\pm$ 0.05 9.54 $\pm$ 0.03
-MetaShift 8.21 $\pm$ 0.09 89.48 $\pm$ 0.28 17.23 $\pm$ 0.34 6.01 $\pm$ 0.30 9.87 $\pm$ 0.05
-OpenImages (Com.) 8.83 92.23 17.44 5.97 9.69
-OpenImages (Rare) 8.60 86.00 17.04 6.19 10.02
-COCO BLIP 12.45 57.11 15.06 8.10 9.10
-NoCaps 12.44 57.61 15.08 8.04 9.28
+Dataset Model $Base$ $Q_{8}$ $ToME$ ${\mathcal{T}}_{intra}^{0.5}$ ${\mathcal{T}}_{inter}^{100}$ ${\mathcal{T}}_{inter}^{1000}$ ${\mathcal{T}}_{inter}^{10000}$
+Waterbirds CLIP 8.48 88.55 17.04 5.84 9.88 9.73 9.91
+CelebA 8.33 89.41 17.25 5.86 9.45 9.54 9.99
+MetaShift 8.21 89.48 17.23 6.01 9.99 9.87 9.75
+OpenImages (Com.) 8.83 92.23 17.44 5.97 9.64 9.69 10.23
+OpenImages (Rare) 8.60 86.00 17.04 6.19 9.69 10.02 9.67
+COCO BLIP 12.45 57.11 15.06 8.10 8.73 9.10 10.31
+NoCaps 12.44 57.61 15.08 8.04 9.49 9.28 10.51
 (b) Runtime (millisecond) $\downarrow$
 Table 2: Wattage and runtime per sample.
-Compared to $Base$, VWTs reduce wattages by up to 19% with ${\mathcal{T}}_{inter}^{1000}$.
+Compared to $Base$, VWTs reduce wattages by up to 25% with ${\mathcal{T}}_{inter}^{100}$.
 Although $Q_{8}$ and $ToME$ display lower wattages than VWTs, they induce significantly longer runtimes, especially the former.
-${\mathcal{T}}_{intra}^{0.5}$ is consistently faster than $Base$, while ${\mathcal{T}}_{inter}^{1000}$ may increase runtimes by up to 20% at most (except on COCO and NoCaps).
+${\mathcal{T}}_{intra}^{0.5}$ is consistently faster than $Base$, while ${\mathcal{T}}_{inter}^{10000}$ may increase runtimes by up to 20% at most (except on COCO and NoCaps).
+Vector Quantization.
+The idea of discretizing continuous distributions has been explored in many works, most recently for image generation.
+Yang et al.
+[68] leveraged clustering for learning a codebook that maps keypoint descriptors to discrete visual words.
+In Wu et al.
+[64] and Bao et al.
+[3], discretization is applied as part of the modelling for ViTs.
+Van Den Oord et al.
+[61] learned discrete image representations by introducing the Vector Quantised-Variational Autoencoder (VQ-VAE) approach.
+Esser et al.
+[15] and Yu et al.
+[71] further improved upon the VQ-VAE by combining the expressiveness of transformers with an adversarial framework.
+Increasingly, vision-language models paired with codebooks conduct image synthesis autoregressively [47, 72, 38, 39, 57, 56, 54].
+Lastly, Yang et al.
+[69] tackled disentangled representation learning and scene decomposition by tokenizing images into separate visual concepts.
+3 Visual Word Tokenizer
+In ViTs, tokenization is a process that splits an image into patches (tokens) which are then projected into a series of embeddings.
+The number of patches is typically fixed (e.g.
+$197$) based on the choice of architecture.
+We seek to split an image into variable length inputs instead (i.e.
+certain images will use $80$ tokens, while others a $100$).
+This variability is induced by conventional text tokenizers [17, 51, 65, 28] for model efficiency.
+We propose to achieve this via visual words that group patches (visual subwords) based on some commonality in the pixel space.
+These visual words capture frequently used patches while retaining infrequent ones as is.
+A simple yet effective grouping can be done using either a criterion that looks at statistics of only one image (an intra-image approach) or across many images (an inter-image approach).
+Figure 2 summarizes the Visual Word Tokenizer (VWT).
+Model Waterbirds CelebA MetaShift OpenImages
+Average $\uparrow$ Worst $\uparrow$ Average $\uparrow$ Worst $\uparrow$ Average $\uparrow$ Worst $\uparrow$ Common $\uparrow$ Rare $\uparrow$
+$Base$ 79.06 21.86 89.61 47.21 95.31 87.69 70.48 63.36
+$Q_{8}$ 79.94 24.25 89.73 48.19 95.23 88.21 70.48 63.19
+$ToME$ 79.80 27.05 89.25 28.52 94.24 87.37 65.69 60.05
+${\mathcal{T}}_{intra}^{0.5}$ 75.56 31.00 89.27 50.93 92.94 86.15 65.52 59.73
+${\mathcal{T}}_{inter}^{100}$ 78.41 26.01 90.04 53.15 90.24 84.28 62.90 58.10
+${\mathcal{T}}_{inter}^{1000}$ 79.19 23.83 90.79 47.96 93.94 86.67 66.15 60.56
+${\mathcal{T}}_{inter}^{10000}$ 79.68 22.90 90.12 46.85 94.81 86.15 69.03 62.50
+(a) CLIP (image classification and subgroup robustness)
+Model COCO NoCaps
+Karpathy $\uparrow$ In-Domain $\uparrow$ Near-Domain $\uparrow$ Out-of-Domain $\uparrow$ Overall $\uparrow$
+BLEU@4 CIDEr CIDEr SPICE CIDEr SPICE CIDEr SPICE CIDEr SPICE
+$Base$ 34.00 107.35 103.57 14.60 98.87 14.11 92.96 13.30 98.34 14.02
+$Q_{8}$ 33.85 106.87 103.86 14.53 100.07 14.18 92.72 13.49 99.12 14.10
+$ToME$ 30.02 94.40 97.35 14.42 90.74 13.39 84.22 12.69 90.37 13.40
+${\mathcal{T}}_{intra}^{0.5}$ 32.80 104.37 99.86 14.20 94.21 13.73 89.52 13.12 94.07 13.68
+${\mathcal{T}}_{inter}^{100}$ 31.10 97.70 95.58 13.78 89.62 13.11 82.42 12.44 89.02 13.08
+${\mathcal{T}}_{inter}^{1000}$ 32.12 101.79 98.19 13.95 93.85 13.49 86.03 12.67 92.89 13.39
+${\mathcal{T}}_{inter}^{10000}$ 33.18 105.08 102.84 14.59 98.03 13.95 91.52 13.14 97.40 13.88
+(b) BLIP (image captioning)
+Table 3: Zero-shot (training-free) image classification (CLIP), subgroup robustness (CLIP) and captioning (BLIP).
+VWTs maintain average accuracy relative to $Base$ while improving robustness on Waterbirds.
+On COCO and NoCaps, VWTs retain a higher performance than $ToME$.
+3.1 An intra-image approach
+The pixel variance of the image patches is the simplest criterion that can be used for grouping.
+In Figure 2, this approach only utilizes the deploy step by grouping the top-k patches with the lowest pixel variance while leaving the rest intact.
+To compress the grouped tokens, we opt to drop them as they tend to exhibit excessive homogeneity.
+We do not include the [CLS] token for dropping.
+This approach is inspired by Minderer et al.
+[42] which aimed to reduce the training cost of OWLv2, an open-vocabulary object detector.
+Dropping patches with the lowest pixel variance removes padding areas and uniform backgrounds from the input mosaics of raw images, thereby increasing training efficiency.
+3.2 An inter-image approach
+Inspired by text tokenizers and codebooks, we propose a variable-length approach that statistically discovers visual words across many images.
+The tokenizer consists of two steps: Pre-Process and Deploy
+Figure 3: Visualization of patch dropping by our visual word tokenizer (${\mathcal{T}}_{intra}^{0.5}$ model).
+Patches with the lowest pixel variance that are dropped are indicated in black.
+In most cases, the dropped patches correspond to the background object on average which is typically uninformative.
+Pre-Process.
+The Bag-of-Visual Words (BoVW) is a popular method for modeling images via discrete representations.
+In Yang et al.
+[68], k-means clustering is applied to keypoint descriptors from SIFT [37] to learn a fixed set of centroids.
+These centroids represent the vocabulary to which multiple descriptors are mapped in a process termed Vector Quantization (VQ).
+In our method, we adopt a variation of this framework by building the BoVW using patches within the pixel space.
+Our design choice is motivated by two main factors.
+First, we find keypoint descriptors to be costly for inference.
+In each forward pass, computing keypoints for each image would significantly increase runtime.
+Second, in our early experimentation, we observed that patches in the embedding space have little similarity to one another.
+Such issues were also described by Bolya et al.
+[5], thus leading to their use of attention scores instead.
+Further justification for leveraging the pixel space is provided in Section 4.6.
+Given a dataset, we first patchify the images using the same patch size as the image encoder (e.g.
+16 for ViT-B/16).
+We then cluster the patches via k-means to form the BoVW of a given vocabulary size, where each centroid represents a visual word.
+Patchification is done via basic tensor operations and not the pre-trained convolutional layer of the ViT to avoid projection into embeddings.
+We also execute this process in batches using the MiniBatchKMeans333from sklearn.cluster import MiniBatchKMeans algorithm due to memory constraints.
+Note that MiniBatchKMeans uses the Euclidean distance by design.
+Since clustering is done in the pixel space, the BoVW may be reused by other ViTs with the same patch size regardless of model type.
 Deploy.
-Once the BoVW is formed, we now turn toward the process of sequence compression.
-One way of leveraging the BoVW would be to merge similar patches in the image space before projecting them into embeddings.
-However, such a naive approach will significantly degrade performance as doing so necessitates the initialization of a new linear layer for projection.
+Once the BoVW is formed, we turn towards the process of sequence compression.
+One way of leveraging the BoVW would be to merge similar patches in the pixel space before projecting them into embeddings.
+However, such a naive approach will significantly degrade performance as the initialization of a new linear layer for projection is required.
 To avoid this, we begin by patchifying and computing the pairwise cosine distance between the patches and BoVW.
@@
 As such, distances higher than a given threshold are masked out to ensure dissimilar patches are not merged.
-This allows us to ensure that infrequently used image patches remain intact.
 At this point, we have determined the groupings of similar patches via their connections to the same visual words.
@@
 4 Experiments
-Consider a pre-trained image encoder $f\in{\mathcal{F}}$ with parameters $\theta\in\Theta$ that transforms inputs $x\in{\mathcal{X}}\subseteq\mathbb{R}^{d}$ to encodings $\hat{x}\in\hat{{\mathcal{X}}}\subseteq\mathbb{R}^{d}$.
+Consider a pre-trained image encoder $f\in{\mathcal{F}}$ with parameters $\theta\in\Theta$ that transforms inputs $x\in{\mathcal{X}}\subseteq\mathbb{R}^{d}$ to encodings $\hat{x}\in\hat{{\mathcal{X}}}\subseteq\mathbb{R}^{\hat{d}}$.
 The encodings can then be mapped to labels $y\in{\mathcal{Y}}$ for some given task, be it classification or generation.
@@
 In our experiments, we seek to analyze the effect of ${\mathcal{T}}$ on online inference efficiency.
-We do so through the lens of (i) performance retention in visual recognition tasks of zero-shot classification, (ii) group robustness, and (iii) generative performance of visual captioning.
-Model Waterbirds CelebA MetaShift OpenImages
-Average $\uparrow$ Worst $\uparrow$ Average $\uparrow$ Worst $\uparrow$ Average $\uparrow$ Worst $\uparrow$ Common $\uparrow$ Rare $\uparrow$
-$Base$ 79.06 $\pm$ 0.14 21.86 $\pm$ 0.32 89.61 $\pm$ 0.14 47.21 $\pm$ 2.41 95.31 $\pm$ 1.00 87.69 $\pm$ 3.08 70.48 63.36
-$Q_{8}$ 79.94 $\pm$ 0.20 24.25 $\pm$ 0.09 89.73 $\pm$ 0.13 48.19 $\pm$ 3.24 95.23 $\pm$ 0.74 88.21 $\pm$ 3.87 70.48 63.19
-$ToME$ 79.80 $\pm$ 0.15 27.05 $\pm$ 0.09 89.25 $\pm$ 0.15 28.52 $\pm$ 4.49 94.24 $\pm$ 0.43 87.37 $\pm$ 3.01 65.69 60.05
-${\mathcal{T}}_{intra}^{0.5}$ 75.56 $\pm$ 0.06 31.00 $\pm$ 0.47 89.27 $\pm$ 0.12 50.93 $\pm$ 1.16 92.94 $\pm$ 0.35 86.15 $\pm$ 2.66 65.52 59.73
-${\mathcal{T}}_{inter}^{1000}$ 79.19 $\pm$ 0.10 23.83 $\pm$ 0.27 90.79 $\pm$ 0.15 47.96 $\pm$ 4.98 93.94 $\pm$ 0.70 86.67 $\pm$ 2.35 66.15 60.56
-(a) CLIP (image classification and group robustness)
-Model COCO NoCaps
-Karpathy $\uparrow$ In-Domain $\uparrow$ Near-Domain $\uparrow$ Out-of-Domain $\uparrow$ Overall $\uparrow$
-BLEU@4 CIDEr CIDEr SPICE CIDEr SPICE CIDEr SPICE CIDEr SPICE
-$Base$ 34.00 107.35 103.57 14.60 98.87 14.11 92.96 13.30 98.34 14.02
-$Q_{8}$ 33.85 106.87 103.86 14.53 100.07 14.18 92.72 13.49 99.12 14.10
-$ToME$ 30.02 94.40 97.35 14.42 90.74 13.39 84.22 12.69 90.37 13.40
-${\mathcal{T}}_{intra}^{0.5}$ 32.80 104.37 99.86 14.20 94.21 13.73 89.52 13.12 94.07 13.68
-${\mathcal{T}}_{inter}^{1000}$ 32.12 101.79 98.19 13.95 93.85 13.49 86.03 12.67 92.89 13.39
-(b) BLIP (image captioning)
-Table 3: Zero-shot image classification (CLIP), group robustness (CLIP) and captioning (BLIP).
-VWTs maintain average accuracy relative to $Base$ while improving robustness on Waterbirds.
-On COCO and NoCaps, VWTs retains a higher performance than $ToME$.
-Dataset and Setting.
-To study the effects of tokenization on (i) performance retention in visual recognition tasks of zero-shot classification, we conduct two sets of evaluations: 1) using one dataset for large-scale evaluation and 2) using three datasets for robustness analysis detailed below.
-We focus on a zero-shot setting by eschewing any end-to-end fine-tuning of $f$.
-For large-scale evaluation, we use the OpenImages v6 dataset [27].
-Following [23], the test split is divided into common and rare subsets that consist of 57 224 and 21 991 images, respectively.
-The former has 214 classes, while the latter has 200.
-To perform classification, we compute the cosine similarity between an image embedding and the encoded text label [^3] The prefix ”a photo of a ” is also added to encode each text label..
-To study performance retention and (ii) group robustness in image classification, we utilize three publicly available datasets: Waterbirds [57], CelebA [33], and MetaShift [29], that are used as benchmarks in robustness and fairness research [47, 32, 64].
-To perform (zero-shot) classification, we compute the cosine similarity between an image embedding and the following encoded text labels3 for Waterbirds, CelebA, and MetaShift, respectively: {’landbird’, ’waterbird’}, {’non-blond’, ’blond’}, {’dog’, ’cat’}.
-The same data splits and spurious attributes as [32] and [64] are used for the former two and latter datasets, respectively.
-Specifically, the following spurious attributes are studied: habitation {’water’, ’land’} for Waterbirds, gender {’female’, ’male’} for CelebA, and context {’indoor’, ’outdoor’} for MetaShift.
-Further details on the subgroups are provided in Section A.1.
-We also compute the variance by shuffling samples among the data splits while maintaining subgroup proportions.
-We leverage attribute annotations not only for analyzing average and robust performance but also for the fairness of tokenization in Section B.4.
-To analyze the effect of ${\mathcal{T}}$ on (iii) generative task of image captioning, we utilize the Karpathy test split of COCO dataset [31] and a validation set of NoCaps dataset  [1] following the setting in previous work [28].
-Finally, to study inference efficiency, we utilize all datasets for the downstream visual tasks (i)-(iii) described above. We focus on the online setting where batch size is 1.
+We focus on the zero-shot setting by eschewing any form of end-to-end fine-tuning for $f$.
 Figure 4: Visualization of the vocabulary of ${\mathcal{T}}_{inter}^{100}$.
@@
 Visual words are shown to depict basic features such as colors or edges.
-As centroids, each visual word is an average representation of patches that belong to the matched cluster.
-Baselines.
+Each visual word is an average representation of patches that belong to the matched cluster.
+4.1 Datasets and Settings
+We conduct our analysis through the lens of (i) classification performance of visual recognition, (ii) subgroup robustness, and (iii) generative performance of visual captioning.
+For (i) and (ii), we utilize three publicly available datasets (Waterbirds [63], CelebA [36], MetaShift [31]) that are typical benchmarks in robustness and fairness research [50, 35, 70].
+To perform (zero-shot) classification, we compute the cosine similarity between an image embedding and the following encoded text labels4 for Waterbirds, CelebA, and MetaShift, respectively: {’landbird’, ’waterbird’}, {’non-blond’, ’blond’}, {’dog’, ’cat’}.
+Further details on the defined subgroups are provided in Appendix 6.1.
+For (i), we also conduct a large-scale evaluation on the OpenImages v6 dataset [29].
+Following Huang et al.
+[25], the test split is divided into common and rare subsets that consist of 57 224 and 21 991 images, respectively.
+The former has 214 classes, while the latter has 200.
+To perform (zero-shot) classification, we compute the cosine similarity between an image embedding and the encoded text label [^4] The prefix ”a photo of a ” is also added to encode each text label..
+For (iii), we utilize the Karpathy test split of COCO dataset [33] and a validation set of NoCaps dataset [1] following the setting in previous work [30].
+Finally, to study inference efficiency, we utilize all datasets for the visual tasks (i)-(iii) described above.
+We compute the efficiency per sample and average across all samples.
+For the wattage calculation, we call ”pynvml.nvmlDeviceGetPowerUsage(handle)/1000” as done by Weights and Biases, which leverages the NVIDIA Management Library for monitoring and managing the NVIDIA GPUs.
+The experiments are also conducted in an isolated environment, thus ensuring the efficiency measurements are as accurate as possible.
+Note that we do not include the Pre-Process step of the inter-image approach into our efficiency calculations as this process is only done once and may be reused multiple times.
+4.2 Implementation Details
 For image classification, we load the pre-trained CLIP [46] model from HuggingFace [^5] https://huggingface.co/openai/clip-vit-base-patch16.
-The prefix ”a photo of a ” is also added to each text category.
 An image size of 224 $\times$ 224 is used with bilinear interpolation for CLIP.
@@
 To perform zero-shot captioning, we use a beam size of 3 along with maximum and minimum lengths of 20 and 5, respectively.
-Unlike [28], we do not include the prefix ”a picture of ” as we found it to work better.
 An image size of 384 $\times$ 384 is used with bicubic interpolation.
@@
 We denote the former as $Q_{8}$ and the latter as $ToME$.
-Following [5], we utilize a reduction per layer for $ToME$ of 13 with CLIP and 23 with BLIP due to their respective input image sizes.
+Following Bolya et al.
+[5], we utilize a reduction per layer for $ToME$ of 13 with CLIP and 23 with BLIP due to their respective input image sizes.
 For the VWTs, we set the top-k of the intra-image approach to 50% of the total number of patches which we denote as ${\mathcal{T}}_{intra}^{0.5}$.
@@
 Lastly, our experiments are conducted using a single NVIDIA A100 GPU.
+Since our focus is on the online setting (real-time), we set the batch size to 1 unless stated otherwise.
 Figure 5: Visualization of long-form captioning on COCO.
-Longer captions are generated via a length penalty of 2.0 and maximum length of 40.
-Interestingly, the smaller vocabulary of ${\mathcal{T}}_{inter}^{100}$ improves descriptiveness and coherence more than ${\mathcal{T}}_{intra}^{0.5}$ or ${\mathcal{T}}_{inter}^{1000}$.
-Inference Efficiency.
+Longer captions are generated via a length penalty of 2.0 and a maximum length of 40.
+Interestingly, the smaller vocabulary of ${\mathcal{T}}_{inter}^{100}$ possesses higher descriptiveness and coherence than ${\mathcal{T}}_{intra}^{0.5}$, ${\mathcal{T}}_{inter}^{1000}$, or ${\mathcal{T}}_{inter}^{10000}$ despite its lesser compression.
+4.3 VWTs and Inference Efficiency
 First we analyze the effects of VWTs on token length.
@@
 The in-domain dataset is represented by the training split if available.
-On text data, in-domain tokenizers [9, 61, 40] have been shown to produce shorter sequences by specializing the vocabulary on the given distribution.
+On text data, in-domain tokenizers [18, 19, 9, 67, 43] have been shown to produce shorter sequences by specializing the vocabulary on the given distribution.
 Interestingly, we observe no such effect with image data as seen by the similar lengths between In-Domain and ImageNet-1K.
-Only on CelebA, do we see a greater reduction with ${\mathcal{T}}_{inter}^{1000}$ and ${\mathcal{T}}_{inter}^{10000}$ pre-processed on ImageNet-1K.
-Second, unlike text tokenizers, a trend of decreasing compression is seen as vocabulary size increases.
+Only on CelebA, do we see a slightly greater reduction with ${\mathcal{T}}_{inter}^{1000}$ and ${\mathcal{T}}_{inter}^{10000}$ pre-processed on ImageNet-1K.
+Second, unlike text tokenizers, decreasing compression is seen as vocabulary size increases.
 With text, larger vocabularies ensure that more tokens are kept as words rather than subwords.
-In our case, we posit that an increasing number of patches are matched to separate visual words, thus lowering the overall compression.
-We restrict our remaining analyses to ${\mathcal{T}}_{inter}^{1000}$ pre-processed on ImageNet-1K unless stated otherwise.
-Having analyzed the effects on token length, we turn our attention to the practical metrics of wattage and runtime.
+We posit that an increasing number of patches are matched to separate visual words, thus lowering the overall compression.
+Having analyzed the effects on token length, we turn to the practical metrics of wattage and runtime.
 For online inference, savings in wattage should not incur a significant cost to runtime.
-In Table 2, we compare the efficiency of tokenization to $Base$, $Q_{8}$, and $ToME$ on the datasets.
+In Table 2, we compare the efficiency of VWTs to $Base$, $Q_{8}$, and $ToME$.
+Note that we compute both metrics using the image encoder of CLIP or BLIP only.
 First, we find the wattage of VWTs to be lower than $Base$ across the datasets.
-On CelebA, this reduction is up to 19% with ${\mathcal{T}}_{inter}^{1000}$.
+On COCO, this reduction is up to 25% with ${\mathcal{T}}_{inter}^{100}$.
 Naturally, efficiency decreases as vocabulary size increases due to smaller compression.
@@
 In particular, $Q_{8}$ results in a significantly longer runtime for $Base$ as tensors need to be repeatedly quantized and dequantized.
-Meanwhile, $ToME$ can increase runtime by up to 2×2$\times$2 × on Waterbirds, CelebA, and MetaShift.
-We observe increases of up to 20% at most with ${\mathcal{T}}_{inter}^{1000}$ on the aforementioned datasets.
-${\mathcal{T}}_{intra}^{0.5}$ displays lower runtimes than $Base$ across the datasets while ${\mathcal{T}}_{inter}^{1000}$ does so on COCO and NoCaps.
+Meanwhile, $ToME$ can increase runtime by up to 100% on Waterbirds, CelebA, MetaShift, and OpenImages.
+We observe increases of up to 20% at most with ${\mathcal{T}}_{inter}^{10000}$ on the aforementioned datasets.
+${\mathcal{T}}_{intra}^{0.5}$ displays lower runtimes than $Base$ across the datasets while ${\mathcal{T}}_{inter}^{100}$, ${\mathcal{T}}_{inter}^{1000}$, and ${\mathcal{T}}_{inter}^{10000}$ do so on COCO and NoCaps.
 We have shown how VWTs are more suitable for online inference by striking a good balance between savings in wattage and costs to runtime.
-Performance Retention of Classification, Robustness, and Captioning.
+Dataset Model ${\mathcal{T}}_{inter}^{100}$ ${\mathcal{T}}_{inter}^{1000}$ ${\mathcal{T}}_{inter}^{10000}$
+In-Domain ImageNet In-Domain ImageNet In-Domain ImageNet
+Waterbirds CLIP 180 179 179 179 182 182
+CelebA 153 154 157 154 170 165
+MetaShift 176 173 175 174 180 180
+OpenImages (Com.) - 171 - 171 - 176
+OpenImages (Rare) - 169 - 169 - 174
+COCO BLIP 405 406 409 408 442 440
+NoCaps - 394 - 396 - 431
+Table 4: Token length per sample (including [CLS]).
+${\mathcal{T}}_{inter}^{{\mathcal{V}}}$ of varying pre-processing data and vocabulary sizes are shown.
+Visual words are formed from patches after the pre-trained convolution layer of CLIP or BLIP.
+Unlike Table 1, poor compression is seen due to dissimilarity between patches and new visual words.
+4.4 VWTs and Visual Performance
 Another important factor in compression is the effect on model performance.
-In Table 3, we tabulate the performance of image classification and captioning using CLIP [43] and BLIP [28], respectively.
-For classification and group robustness, we analyze the average and worst-group (robust) performance as done by [47, 46].
-First, we find the degradation in average accuracy to be small for ${\mathcal{T}}_{intra}^{0.5}$ and ${\mathcal{T}}_{inter}^{1000}$.
-The largest drop of up to 2% only is shown by the former on MetaShift.
-Likewise, $Q_{8}$ and $ToME$ retain a high average accuracy across the datasets.
-Second, we observe possible improvements in robustness with VWTs.
-On Waterbirds and CelebA, the worst-group accuracy with ${\mathcal{T}}_{intra}^{0.5}$ increases by up to 29% and 8%, respectively.
-Only on MetaShift is robustness lower than $Q_{8}$ or $ToME$ relative to $Base$.
-Like [20], we find compression to not always be harmful to group robustness.
-We conduct more evaluations on the large-scale OpenImages v6 dataset for zero-shot classification.
-We report mean Average Precision (mAP) as done by [23].
-From Tables 2 and 3, we find VWTs to be more suitable for online inference than $Q_{8}$ or $ToME$ in the large-scale setting, while retaining comparable performance to the $Base$ model.
-For image captioning with BLIP, we evaluate our models following the setting in  [28] by using the BLEU, CIDEr, and SPICE metrics w.r.t.
+In Table 3, we tabulate the performance of image classification and captioning using CLIP and BLIP, respectively.
+For visual recognition and subgroup robustness, we analyze performance from an average and worst-group perspective as done by Sagawa et al.
+[50] and Romiti et al.
+[49].
+First, we find the degradation in average accuracy to be small for the VWTs.
+The largest drop of up to 2% only is shown by ${\mathcal{T}}_{intra}^{0.5}$ on MetaShift.
+Likewise, both $Q_{8}$ and $ToME$ retain a high average accuracy across the datasets.
+Second, we observe possible improvements in subgroup robustness with VWTs.
+On Waterbirds and CelebA, the worst-group accuracy (WGA) with ${\mathcal{T}}_{intra}^{0.5}$ increases by up to 29% and 8%, respectively.
+Only on MetaShift do we observe a lower WGA than $Q_{8}$ or $ToME$ relative to $Base$.
+Like Gee et al.
+[20], we find compression to not always be harmful to subgroup robustness by improving the WGA at a negligible cost to overall performance.
+To further validate performance, we conduct additional evaluations on the large-scale OpenImages v6 dataset for zero-shot classification.
+We report mean Average Precision (mAP) as done by Huang et al.
+[25].
+Likewise, from Tables 2 and 3, we find the VWTs to be more suitable for online inference than $Q_{8}$ or $ToME$ in the large-scale setting while retaining comparable performance to $Base$.
+For image captioning with BLIP, we evaluate our models following the setting in Li et al.
+[30] by using the BLEU, CIDEr, and SPICE metrics w.r.t.
 the ground truth captions.
-On COCO, VWTs display a higher performance than $ToME$ and are competitive with $Q_{8}$.
-On NoCaps, we see the largest degradation on the out-of-domain samples with ${\mathcal{T}}_{inter}^{1000}$.
-However, overall performance is still higher than $ToME$.
+On COCO, the VWTs display a higher performance than $ToME$ and are competitive with $Q_{8}$.
+On NoCaps, we see the largest degradation on the out-of-domain samples with ${\mathcal{T}}_{inter}^{100}$.
+However, overall performance of ${\mathcal{T}}_{intra}^{0.5}$, ${\mathcal{T}}_{inter}^{1000}$, and ${\mathcal{T}}_{inter}^{10000}$ are still higher than $ToME$.
 Only $Q_{8}$ displays a slight improvement over $Base$.
-Like $ToME$, VWTs are shown to not require additional end-to-end fine-tuning for performance retention.
-(a) Wattage (watt) $\downarrow$
-(b) Runtime (millisecond) $\downarrow$
-[…]
-Visual Words.
-To better understand the inter-image approach, we visualize the patch matching of ${\mathcal{T}}_{inter}^{100}$ in Figure 1.
-We highlight in identical colors patches that are matched with one another.
-First, we find that patches representing the background are more frequently grouped than those of the foreground.
-On Waterbirds, the merging of background patches may explain the improved robustness in Table 3 by mitigating the spurious correlation with the background.
-On CelebA, ${\mathcal{T}}_{inter}^{100}$ tends to avoid matching the eyes or mouths of the individuals.
-Second, we observe that patch matching is capable of grouping similar but non-adjacent visual concepts.
-In certain examples of MetaShift and NoCaps, multiple cats and foods are seemingly matched together, respectively.
-Our analysis shows that patch matching serves as a rudimentary form of image segmentation.
-In Figure 3, we visualize the vocabulary of ${\mathcal{T}}_{inter}^{100}$ to analyze the formation of the visual words.
-We show patches from ImageNet-1K (i.e.
-pre-processing data) that are matched to each visual word using the Euclidean distance.
-Since visual words are centroids, patches that are matched to the same visual word belong to the same cluster.
-We observe that visual words depict basic features such as colors or edges.
-These features are also formed as an average representation of the patches that belong to each cluster.
-By representing basic features, visual words serve as effective anchor points to which patches can be matched to.
-Analysis on the sparsity of the vocabularies in the inter-image approach can be found in Section B.5.
-Long-form Captioning.
+Like $ToME$, the VWTs are shown to not require additional end-to-end fine-tuning for performance retention.
 The generated captions by BLIP [30] can be used as priors for further training of other models.
@@
 With longer captions, the generation may degenerate into unnecessary repetitions on certain samples.
-Interestingly, descriptiveness and coherence improves more with ${\mathcal{T}}_{inter}^{100}$ than ${\mathcal{T}}_{intra}^{0.5}$ or ${\mathcal{T}}_{inter}^{1000}$.
-Additional long-form captions by $Q_{8}$, $ToME$, and ${\mathcal{T}}_{inter}^{10000}$ are provided in Section C.1.
-Batch Processing.
-Earlier we stated how VWTs works optimally for online inference.
-In Figure 5, we seek to better understand its effectiveness for offline inference where batch sizes are greater than 1.
-Using typical batches of $\{4,8,16,32\}$, we compare ${\mathcal{T}}_{intra}^{0.5}$ and ${\mathcal{T}}_{inter}^{1000}$ to $Base$ and $ToME$.
-First, we find that as batch size increases, lower wattages relative to $Base$ are seen with ${\mathcal{T}}_{intra}^{0.5}$ and $ToME$, particularly at lower batch sizes.
-Both VWTs and $ToME$ eventually reach wattage parity with $Base$ at larger batches (i.e.
-32), with ${\mathcal{T}}_{inter}^{1000}$ doing so much earlier.
-Second, we observe different runtime trends between VWTs and $ToME$.
-As batches grow, ${\mathcal{T}}_{inter}^{1000}$ continues to show similar runtimes to $Base$.
-Meanwhile, ${\mathcal{T}}_{intra}^{0.5}$ and $ToME$ display lower runtimes with the exception of batch size 4 for the latter.
-As such, during batching, the inter-image approach does not confer the same degree of efficiency improvements.
-When to use the intra-image or inter-image approach?
-Concerning the choice between the intra-image or inter-image approach, we posit that when global information is required, dropping tokens may be more advantageous by removing unnecessary noise from the visual input.
-On the other hand, when the task necessitates local information (e.g.
-long-form captioning, object detection, etc.), merging tokens may better preserve the visual concepts.
-For example, the image on the first row of Figure 7 shows that ${\mathcal{T}}_{intra}^{0.5}$ removes the mountainous background, thus leading to the absence of ”mountain” in the long-form caption of Figure 4.
-Potentially, combining both approaches may maximize compression while preserving information by merging the mountains and dropping the bushes in said image.
-5 Conclusion
-In this work, we set out to define a training-free tokenization for ViTs that lowers wattage while balancing costs to runtime and performance.
-In online scenarios, we have shown empirically that our intra-image and inter-image approaches are competitive with 8-bit quantization and token merging for image classification and captioning.
-Qualitatively, we observe how the inter-image approach groups analogous visual concepts via visual words that represent basic features.
-Analysis on large-scale classification further validates the viability of our method while long-form captioning shows its potential for improving descriptiveness and coherence.
-Lastly, the inter-image approach has limits for batch processing due to its non-parallelism.
-As a future work, further research can be made into combining both approaches.
-Acknowledgements
-This research was supported by a European Research Council (ERC) Starting Grant for the project “Bayesian Models and Algorithms for Fairness and Transparency”, funded under the European Union’s Horizon 2020 Framework Programme (grant agreement no.
-851538).
-Novi Quadrianto is also supported by the Basque Government through the BERC 2022-2025 program and by the Ministry of Science and Innovation: BCAM Severo Ochoa accreditation CEX2021-001142-S / MICIN/ AEI/ 10.13039/501100011033.
-Viktoriia Sharmanska is currently at Epic Games.
-Appendix A Further Details
-A.1 Datasets for Group Robustness
-Here, we detail the task of each dataset for group robustness. In Table 4, we also tabulate the labels and attributes that define each subgroup along with their sample sizes.
-Waterbirds.
-Given an image of a bird, the task is to predict whether it is a waterbird or landbird [57]. The attribute is the background that the bird is on [47].
-CelebA.
-Given an image of a person, the task is to predict whether their hair color is blond or not [33]. The attribute is the binary gender of the person.
-MetaShift.
-Given an image of an animal, the task is to predict whether it is a dog or cat [29]. The attribute is the environment that the dog or cat is in.
-Dataset Subgroup Label Attribute Training Validation Test
-Waterbirds 0 0 (landbird) 0 (on land) 3498 467 2255
-1 0 (landbird) 1 (on water) 184 466 2255
-2 1 (waterbird) 0 (on land) 56 133 642
-3 1 (waterbird) 1 (on water) 1057 133 642
-CelebA 0 0 (non-blond) 0 (woman) 71629 8535 9767
-1 0 (non-blond) 1 (man) 66874 8276 7535
-2 1 (blond) 0 (woman) 22880 2874 2480
-3 1 (blond) 1 (man) 1387 182 180
-MetaShift 0 0 (dog) 0 (outdoor) 784 127 273
-1 0 (dog) 1 (indoor) 507 75 191
-2 1 (cat) 0 (outdoor) 196 33 65
-3 1 (cat) 1 (indoor) 789 114 345
-Table 4: Defined subgroups in Waterbirds, CelebA, and MetaShift.
-Appendix B Supplementary Experiments
-Here, we detail the supplementary experiments that we conducted.
-First, we explore additional design choices and effects of the inter-image approach.
-Section B.1 shows the ineffectiveness of patch matching with visual words formed in the embedding space.
-In Section B.2, random matching is applied to further determine if matching by the visual words are meaningful.
-Section B.3 ablates the similarity threshold by relaxing it for further compression.
-The fairness of tokenization and sparsity of the vocabulary are shown in Sections B.4 and B.5, respectively.
-B.1 Ineffectiveness of the Embedding Space
-In Table 5, we analyze the sequence compression using visual words formed in the embedding space.
-To do so, we initialize each vocabulary by clustering embeddings from the pre-trained convolution layer of CLIP or BLIP.
-During inference, we match the activations from the pre-trained convolution layer with the visual words.
-Compared to Table 1, we observe a notable reduction in the degree of compression irrespective of pre-processing data and vocabulary size.
-In [50], embeddings are shown to become progressively more similar to one another due to self-attention.
-Hence, it is unsurprising that matching activations and visual words at the beginning of the network is ineffective.
-B.2 Random Matching of Visual Words
-In Table 6, we study the effects of randomly matching the patches.
-We initialize the pairwise cosine distance by sampling from a uniform distribution of $[0,2]$.
-First, we find the token length (including [CLS]) to differ noticeably from Table 1.
-For ${\mathcal{T}}_{inter}^{100}$, sequences are further reduced across the datasets.
-Conversely, ${\mathcal{T}}_{inter}^{1000}$ and ${\mathcal{T}}_{inter}^{10000}$ display little compression.
-Second, performance is shown to change from Table 3.
-We observe a significant degradation in average and worst-group accuracies with ${\mathcal{T}}_{inter}^{100}$ on Waterbirds and MetaShift.
-For ${\mathcal{T}}_{inter}^{1000}$ and ${\mathcal{T}}_{inter}^{10000}$, performance does not shift much from $Base$ as barely any compression occurs.
-With random matching, the captions become completely different than those of $Base$ in Figure 10.
-B.3 Ablation of the Similarity Threshold
-We have shown how VWTs can improve the efficiency of online inference.
-To better understand their limitations, we ablate the threshold of ${\mathcal{T}}_{inter}^{100}$ by setting it to values of $\{0.2,0.3,0.4,0.5\}$ in Table 7.
-We seek to determine if exploiting higher thresholds for increased compression is a viable strategy.
-Naturally, we observe a reduction in performance as increasingly dissimilar patches are merged.
-For Waterbirds and MetaShift, the worst-group accuracy degrades more significantly than the average, especially with the former.
-Interestingly, average accuracy remains relatively unchanged while worst-group accuracy improves significantly on CelebA irrespective of similarity threshold.
-We posit that at higher thresholds, the merging of core features represented by the foreground object results in the reduced performance of Waterbirds and MetaShift.
-B.4 Fairness of the Tokenization
-On text data, tokenizers have been shown to induce unfairness between languages [41, 2].
-Since most LLMs are built using English data, tokenization of minority languages produces inherently longer sequences that raises compute costs.
-We seek to analyze if similar effects exist with VWTs as well.
-In Table 8, we show the breakdown in token length (including [CLS]) and accuracy (w.r.t $Base$) by subgroup.
-First, we observe a notable difference in compression between the subgroups of Waterbirds.
-With ${\mathcal{T}}_{100}$, sequences might differ by up to 39 tokens as seen with subgroups 0 and 3.
-Smaller discrepancies are displayed on CelebA and MetaShift except for ${\mathcal{T}}_{100}$ on the former.
-Second, we find that compression does not affect all subgroups equally.
-Accuracy improves on certain subgroups and degrades on others.
-A stronger compression does not correlate also with a large change in performance.
-Like text tokenizers, VWTs may induce inequality in compression that leads to varying inference costs and performance for different subgroups.
-B.5 Sparsity of the Vocabulary
-To better understand the utilization of the visual words, we plot the probability distribution of the matches in Figure 6.
-Regardless of the dataset, we find that certain visual words are matched more frequently than others, thus leading to a large skew in the distributions.
-Greater sparsity is also displayed by larger vocabularies as many visual words remain unused across datasets.
-As such, pruning may be applied to achieve a more efficient vocabulary size post-clustering.
-Appendix C Additional Results
-Here, we provide additional results by other models and hyperparameters in our experiments.
-Section C.1 lists long-form captions by $Q_{8}$, $ToME$, and ${\mathcal{T}}_{inter}^{10000}$.
-Sections C.2 and C.3 detail the intra-image and inter-image approaches with various dropping ratios and vocabulary sizes, respectively.
-C.1 Long-form Captions
-In Figure 7, we showcase long-form captions by $Q_{8}$, $ToME$, and ${\mathcal{T}}_{inter}^{10000}$.
-First, we observe improvements in the accuracy of the descriptions with $Q_{8}$ over $Base$ in Figure 4.
-Although the captions relate better to the image, the overall length is shorter than those of other methods.
-The generation may continue to introduce unnecessary repetitions.
-Second, we find $ToME$ and ${\mathcal{T}}_{inter}^{10000}$ to also be beneficial for descriptiveness and coherence.
-On certain samples, repeating n-grams may still occur.
-C.2 Dropping Ratio
-Following [39], we apply the intra-image approach with ratios of $\{0.25,0.33,0.7\}$ to the total number of patches in Table 9.
-We observe a natural degradation as the ratio increases with performance dropping steeply at 0.7.
-In Figures 8 and 9, we visualize the dropped patches and captions on COCO, respectively.
-In most cases, the patches with the lowest pixel variance correspond to uninformative backgrounds.
-Dropping such patches may continue to preserve the foreground object.
-C.3 Vocabulary Size
-In Table 10, we apply the inter-image approach with vocabulary sizes of $\{100,10000\}$.
-We find performance to be noticeably lower than $Base$ in Table 3 with the smaller vocabulary of ${\mathcal{T}}_{inter}^{100}$.
-Meanwhile, performance retention is higher for ${\mathcal{T}}_{inter}^{10000}$ due to lower sequence compression in Table 1.
-When visualizing the captions on COCO in Figure 10, we also observe that increasing thresholds cause larger deviations from $Base$ than reducing the vocabulary size.
-Dataset ${\mathcal{T}}_{inter}^{100}$ ${\mathcal{T}}_{inter}^{1000}$ ${\mathcal{T}}_{inter}^{10000}$
-In-Domain ImageNet In-Domain ImageNet In-Domain ImageNet
-Waterbirds 180 179 179 179 182 182
-CelebA 153 154 157 154 170 165
-MetaShift 176 173 175 174 180 180
-COCO 405 406 409 408 442 440
-NoCaps - 394 - 396 - 431
-Table 5: Token length per sample (including [CLS]).
-Tokenizers of varying pre-processing data and vocabulary sizes are shown.
-Visual words are formed from embeddings in the pre-trained convolution layer of CLIP or BLIP and not the image patches.
-Poor compression is seen irrespective of pre-processing data and vocabulary size due to the lack of similarity between the activations and visual words.
-Dataset Model ${\mathcal{T}}_{intra}^{100}$ ${\mathcal{T}}_{intra}^{1000}$ ${\mathcal{T}}_{intra}^{10000}$
+Interestingly, descriptiveness and coherence improves more with ${\mathcal{T}}_{inter}^{100}$ than ${\mathcal{T}}_{intra}^{0.5}$, ${\mathcal{T}}_{inter}^{1000}$, or ${\mathcal{T}}_{inter}^{10000}$ inspite of its higher sequence compression as seen on COCO in Table 1.
+Dataset Model ${\mathcal{T}}_{inter}^{100}$ ${\mathcal{T}}_{inter}^{1000}$ ${\mathcal{T}}_{inter}^{10000}$
 Waterbirds CLIP 88 179 195
@@
 Average $\uparrow$ Worst $\uparrow$ Average $\uparrow$ Worst $\uparrow$ Average $\uparrow$ Worst $\uparrow$ Common $\uparrow$ Rare $\uparrow$
-${\mathcal{T}}_{inter}^{100}$ 66.60 $\pm$ 0.28 9.61 $\pm$ 1.51 90.24 $\pm$ 0.06 56.48 $\pm$ 4.32 76.16 $\pm$ 0.70 61.43 $\pm$ 1.32 41.06 39.84
-${\mathcal{T}}_{inter}^{1000}$ 78.10 $\pm$ 0.18 22.74 $\pm$ 1.24 90.25 $\pm$ 0.15 47.96 $\pm$ 2.24 94.20 $\pm$ 0.63 87.18 $\pm$ 1.78 70.29 63.48
-${\mathcal{T}}_{inter}^{10000}$ 78.98 $\pm$ 0.07 21.96 $\pm$ 0.31 89.61 $\pm$ 0.07 47.46 $\pm$ 2.62 95.27 $\pm$ 0.97 87.18 $\pm$ 2.35 70.52 63.40
-(b) CLIP (image classification and group robustness)
+${\mathcal{T}}_{inter}^{100}$ 66.60 9.61 90.24 56.48 76.16 61.43 41.06 39.84
+${\mathcal{T}}_{inter}^{1000}$ 78.10 22.74 90.25 47.96 94.20 87.18 70.29 63.48
+${\mathcal{T}}_{inter}^{10000}$ 78.98 21.96 89.61 47.46 95.27 87.18 70.52 63.40
+(b) CLIP (image classification and subgroup robustness)
 Model COCO NoCaps
@@
 (c) BLIP (image captioning)
-Table 6: Token length (including [CLS]) and performance of VWTs with random matching.
+Table 5: Token length (including [CLS]) and performance of the inter-image approach with random merging.
 The pairwise cosine distance is initialized by sampling from a uniform distribution of $[0,2]$.
-Average and worst-group accuracies degrade noticeably with ${\mathcal{T}}_{100}$ except for CelebA.
-Performance does not change much for ${\mathcal{T}}_{1000}$ and ${\mathcal{T}}_{10000}$ from $Base$ since barely any compression occurs.
+Average and worst-group accuracies degrade noticeably with ${\mathcal{T}}_{inter}^{100}$ except for CelebA.
+Performance does not change much for ${\mathcal{T}}_{inter}^{1000}$ and ${\mathcal{T}}_{inter}^{10000}$ from $Base$ since barely any compression occurs.
+4.5 Visualization of the Visual Words
+To better understand the VWT, we visualize the patch matching of ${\mathcal{T}}_{intra}^{0.5}$ and ${\mathcal{T}}_{inter}^{100}$ in Figures 3 and 1, respectively.
+For the latter, we highlight in identical colors patches that are matched with one another.
+With ${\mathcal{T}}_{intra}^{0.5}$, the patches with the lowest pixel variance typically correspond to uninformative backgrounds.
+In most cases, the dropping of such patches continues to preserve the foreground object.
+With ${\mathcal{T}}_{inter}^{100}$, we find that patches representing the background are more frequently grouped than those of the foreground object.
+On Waterbirds, the merging of background patches may explain the improved robustness in Table 3 by mitigating the spurious correlation with the background.
+On CelebA, ${\mathcal{T}}_{inter}^{100}$ tends to avoid matching the eyes or mouths of the individuals.
+We also observe that patch matching is capable of grouping similar but non-adjacent visual concepts.
+In certain examples of MetaShift and NoCaps, multiple cats and foods are seemingly matched together, respectively.
+Our analysis shows that patch matching serves as a rudimentary form of image segmentation by identifying similar visual concepts for compression.
+In Figure 4, we visualize the vocabulary of ${\mathcal{T}}_{inter}^{100}$ to analyze the formation of the visual words.
+We show patches from ImageNet-1K (i.e.
+pre-processing data) that are matched to each visual word using the Euclidean distance.
+Since visual words are centroids, patches that are matched to the same visual word belong to the same cluster.
+We observe that visual words depict basic features such as colors or edges.
+These features are also formed as an average representation of the patches that belong to each cluster.
+By representing basic features, visual words serve as effective anchor points to which patches can be matched to.
 Dataset Thresh Length Average $\uparrow$ Worst $\uparrow$
-Waterbirds 0.2 97 76.71 $\pm$ 0.24 24.09 $\pm$ 0.59
-0.3 79 74.85 $\pm$ 0.15 21.13 $\pm$ 0.36
-0.4 66 73.35 $\pm$ 0.12 16.30 $\pm$ 0.24
-0.5 58 72.81 $\pm$ 0.10 13.14 $\pm$ 0.48
-CelebA 0.2 66 89.44 $\pm$ 0.08 56.48 $\pm$ 1.70
-0.3 55 88.85 $\pm$ 0.26 58.70 $\pm$ 4.21
-0.4 50 88.31 $\pm$ 0.14 57.04 $\pm$ 3.26
-0.5 48 88.21 $\pm$ 0.32 55.74 $\pm$ 1.40
-MetaShift 0.2 84 86.16 $\pm$ 0.89 77.46 $\pm$ 2.34
-0.3 69 81.16 $\pm$ 1.00 67.77 $\pm$ 1.43
-0.4 60 77.73 $\pm$ 1.03 60.65 $\pm$ 2.57
-0.5 55 74.75 $\pm$ 1.36 56.08 $\pm$ 3.38
+Waterbirds 0.2 97 76.71 24.09
+0.3 79 74.85 21.13
+0.4 66 73.35 16.30
+0.5 58 72.81 13.14
+CelebA 0.2 66 89.44 56.48
+0.3 55 88.85 58.70
+0.4 50 88.31 57.04
+0.5 48 88.21 55.74
+MetaShift 0.2 84 86.16 77.46
+0.3 69 81.16 67.77
+0.4 60 77.73 60.65
+0.5 55 74.75 56.08
 Table 6: Token length (including [CLS]) and performance using ${\mathcal{T}}_{inter}^{100}$ with varying similarity thresholds.
 Higher thresholds can reduce sequences to 48 tokens on CelebA.
-Greater compression results in higher degradation in performance, particularly in robustness.
-Model Subgroup Waterbirds CelebA MetaShift
-Length $\Delta$ Accuracy Length $\Delta$ Accuracy Length $\Delta$ Accuracy
-$Base$ 0 197 98.89 197 95.63 197 98.54
-1 82.45 96.23 93.19
-2 21.86 48.67 87.69
-3 54.73 50.00 95.36
-${\mathcal{T}}_{inter}^{100}$ 0 144 -0.67 85 -3.25 118 -0.74
-1 106 -3.82 88 -0.13 110 -6.18
-2 140 19.00 98 32.62 119 -2.92
-3 105 6.17 99 6.30 100 -9.02
-${\mathcal{T}}_{inter}^{1000}$ 0 160 -0.40 119 -0.46 138 -0.25
-1 129 -0.22 117 0.74 138 -2.06
-2 160 9.03 124 18.92 143 -1.17
-3 129 2.18 124 -4.07 131 -2.13
-${\mathcal{T}}_{inter}^{10000}$ 0 180 -0.09 158 0.25 163 -0.37
-1 156 1.34 151 0.66 167 0.19
-2 181 4.75 158 2.96 169 -1.75
-3 158 1.71 156 -6.30 163 -0.81
-Table 8: Distribution of token length (including [CLS]) and accuracy (w.r.t.
-$Base$) by subgroup.
-Tokenizers of varying vocabulary sizes are shown.
-Like text tokenizers, VWTs may induce significantly unequal token lengths as seen with ${\mathcal{T}}_{inter}^{100}$ on Waterbirds.
-Subgroups are also affected unequally as stronger compression does not correlate with greater improvement or degradation in performance.
-(a) Waterbirds
-(b) CelebA
-(c) MetaShift
-Figure 6: Probability distribution of the matched visual words.
-Tokenizers of varying vocabulary sizes are shown.
-The probability distribution exhibits a large skew irrespective of the dataset as certain visual words are matched more frequently than others.
-Larger vocabularies display greater sparsity as the many visual words that remain unmatched may be pruned.
-Figure 7: Visualization of long-form captions by $Q_{8}$, $ToME$, and ${\mathcal{T}}_{inter}^{10000}$ on COCO.
-Dataset Model ${\mathcal{T}}_{intra}^{0.25}$ ${\mathcal{T}}_{intra}^{0.33}$ ${\mathcal{T}}_{intra}^{0.7}$
-Waterbirds CLIP 148 132 59
-CelebA
-MetaShift
-OpenImages (Com.)
-OpenImages (Rare)
-COCO BLIP 433 386 173
-NoCaps
-(a) Length
-Model Waterbirds CelebA MetaShift OpenImages
-Average $\uparrow$ Worst $\uparrow$ Average $\uparrow$ Worst $\uparrow$ Average $\uparrow$ Worst $\uparrow$ Common $\uparrow$ Rare $\uparrow$
-${\mathcal{T}}_{intra}^{0.25}$ 78.31 $\pm$ 0.09 26.69 $\pm$ 0.24 89.46 $\pm$ 0.16 48.47 $\pm$ 1.56 94.70 $\pm$ 0.43 87.69 $\pm$ 1.54 69.79 63.37
-${\mathcal{T}}_{intra}^{0.33}$ 77.72 $\pm$ 0.24 27.78 $\pm$ 0.48 89.31 $\pm$ 0.22 48.83 $\pm$ 1.42 94.62 $\pm$ 0.41 87.18 $\pm$ 2.35 69.11 62.75
-${\mathcal{T}}_{intra}^{0.7}$ 73.27 $\pm$ 0.12 18.80 $\pm$ 0.32 89.19 $\pm$ 0.05 45.37 $\pm$ 1.28 82.61 $\pm$ 1.10 63.59 $\pm$ 0.89 46.83 42.68
-(b) CLIP (image classification and group robustness)
-Model COCO NoCaps
-Karpathy $\uparrow$ In-Domain $\uparrow$ Near-Domain $\uparrow$ Out-of-Domain $\uparrow$ Overall $\uparrow$
-BLEU@4 CIDEr CIDEr SPICE CIDEr SPICE CIDEr SPICE CIDEr SPICE
-${\mathcal{T}}_{intra}^{0.25}$ 33.92 107.47 103.59 14.65 98.39 14.07 94.47 13.59 98.34 14.06
-${\mathcal{T}}_{intra}^{0.33}$ 33.59 106.25 103.19 14.65 97.40 13.95 92.66 13.41 97.27 13.95
-${\mathcal{T}}_{intra}^{0.7}$ 29.64 93.20 89.89 13.53 85.04 12.82 80.81 12.31 84.88 12.82
-(c) BLIP (image captioning)
-Table 9: Token length (including [CLS]) and performance of VWTs with varying dropping ratios.
-Figure 8: Visualization of patch dropping by our visual word tokenizer (${\mathcal{T}}_{intra}^{0.5}$ model).
-Patches with the lowest pixel variance that are dropped are indicated in black.
-In most cases, the dropped patches correspond to the background object that is typically uninformative.
-Figure 9: Visualization of captions by token dropping on COCO.
-Token dropping of up to 0.5 may not cause significant deviation in the generations from the base model.
-On certain samples, the accuracy of the descriptions can improve as well.
-Model Waterbirds CelebA MetaShift OpenImages
-Average $\uparrow$ Worst $\uparrow$ Average $\uparrow$ Worst $\uparrow$ Average $\uparrow$ Worst $\uparrow$ Common $\uparrow$ Rare $\uparrow$
-${\mathcal{T}}_{inter}^{100}$ 78.41 $\pm$ 0.09 26.01 $\pm$ 0.41 90.04 $\pm$ 0.22 53.15 $\pm$ 3.21 90.24 $\pm$ 0.40 84.28 $\pm$ 2.38 62.90 58.10
-${\mathcal{T}}_{inter}^{10000}$ 79.68 $\pm$ 0.11 22.90 $\pm$ 0.16 90.12 $\pm$ 0.18 46.85 $\pm$ 4.63 94.81 $\pm$ 0.95 86.15 $\pm$ 1.54 69.03 62.50
-(a) CLIP (image classification and group robustness)
-Model COCO NoCaps
-Karpathy $\uparrow$ In-Domain $\uparrow$ Near-Domain $\uparrow$ Out-of-Domain $\uparrow$ Overall $\uparrow$
-BLEU@4 CIDEr CIDEr SPICE CIDEr SPICE CIDEr SPICE CIDEr SPICE
-${\mathcal{T}}_{inter}^{100}$ 31.10 97.70 95.58 13.78 89.62 13.11 82.42 12.44 89.02 13.08
-${\mathcal{T}}_{inter}^{10000}$ 33.18 105.08 102.84 14.59 98.03 13.95 91.52 13.14 97.40 13.88
-(b) BLIP (image captioning)
-Table 10: Token length (including [CLS]) and performance of VWTs with varying vocabulary sizes.
-Figure 10: Visualization of captions by VWTs on COCO.
-Increasing the similarity threshold causes the generated captions to deviate more than reducing the vocabulary size.
+Greater compression results in higher degradation in performance, particularly in subgroup robustness.
+4.6 Ineffectiveness of the Embedding Space
+In Table 4, we analyze the sequence compression using visual words formed in the embedding space.
+Instead of initializing the vocabulary by clustering patches in the pixel space, we do so with patches after the pre-trained convolution layer of CLIP or BLIP.
+During inference, we match the patches after the pre-trained convolution layer with these new visual words.
+Compared to Table 1, we observe a notable reduction in the compression irrespective of pre-processing data and vocabulary size.
+In Tang et al.
+[55], embeddings are shown to become progressively more similar to one another due to self-attention.
+As such, it is unsurprising that matching patches and visual words in the initial embedding space is found to be ineffective.
+4.7 Random Merging of Tokens
+In Table 5, we study the effects of randomly merging the tokens.
+We initialize the pairwise cosine distance by sampling from a uniform distribution of $[0,2]$.
+First, we find the token length (including [CLS]) to differ noticeably from Table 1.
+For ${\mathcal{T}}_{inter}^{100}$, sequences are further reduced across the datasets.
+Conversely, ${\mathcal{T}}_{inter}^{1000}$ and ${\mathcal{T}}_{inter}^{10000}$ display little compression.
+Second, performance is shown to change from Table 3.
+We observe a significant degradation in average and worst-group accuracies with ${\mathcal{T}}_{inter}^{100}$ on Waterbirds and MetaShift.
+For ${\mathcal{T}}_{inter}^{1000}$ and ${\mathcal{T}}_{inter}^{10000}$, performance does not shift much from $Base$ as barely any compression occurs.
+With random merging, the captions are shown to deviate completely from those of $Base$ in Figure 6, thus further demonstrating that visual words are an effective criterion for grouping patches as shown in Figure 1.
+4.8 Ablation of the Similarity Threshold
+We have shown how the inter-image approach can improve the efficiency of online inference.
+To better understand their limitations, we ablate the threshold of ${\mathcal{T}}_{inter}^{100}$ by setting it to values of $\{0.2,0.3,0.4,0.5\}$ in Table 6.
+We seek to determine if exploiting higher thresholds for increased compression is a viable strategy.
+Naturally, we observe a reduction in performance as increasingly dissimilar patches are merged.
+For Waterbirds and MetaShift, the WGA degrades more significantly than the average, especially with the former.
+Interestingly, average accuracy remains relatively unchanged while WGA improves significantly on CelebA irrespective of similarity threshold.
+We posit that at higher thresholds, the merging of core features represented by the foreground object results in the reduced performance of Waterbirds and MetaShift.
+Figure 6: Visualization of image captions on COCO by the inter-image approach.
+The generated captions are shown to deviate more when increasing the similarity threshold than when reducing the vocabulary size.
 With random matching, the model begins to completely misunderstand the image.
+4.9 When to use the intra-image or inter-image approach?
+Concerning the choice between the intra-image or inter-image approach, we posit that when global information is required, dropping tokens may be more advantageous by removing unnecessary noise from the visual input.
+On the other hand, when the task necessitates local information (e.g.
+long-form captioning), merging tokens may better preserve the visual concepts.
+For example, the image on the first row of Figure 7 of Appendix 6.2 shows that ${\mathcal{T}}_{intra}^{0.5}$ removes the mountainous background, thus leading to the absence of ”mountain” in the long-form caption of Figure 5.
+Potentially, combining both approaches may maximize compression while preserving information by merging the mountains and dropping the bushes.
+5 Conclusion
+In this work, we set out to define a training-free tokenization for ViTs that lowers wattage while balancing costs to runtime and performance.
+In online scenarios, we have shown empirically that our intra-image and inter-image approaches are stronger than 8-bit quantization and token merging for image classification and captioning.
+Analysis on large-scale classification further validates the viability of our method while long-form captioning shows its potential for improving descriptiveness and coherence.
+Qualitatively, we observe how the criterion of the intra-image approach typically corresponds to the background while that of the inter-image approach groups analogous visual concepts based on visual words that represent basic features.
+As a future work, additional research could explore combining both the intra-image and inter-image approaches, potentially improving further the performance and efficiency of the VWT.
+Acknowledgements
+This research was supported by a European Research Council (ERC) Starting Grant for the project “Bayesian Models and Algorithms for Fairness and Transparency”, funded under the European Union’s Horizon 2020 Framework Programme (grant agreement no.
+851538).
+Novi Quadrianto is also supported by the Basque Government through the BERC 2022-2025 program and by the Ministry of Science and Innovation: BCAM Severo Ochoa accreditation CEX2021-001142-S / MICIN/ AEI/ 10.13039/501100011033.
+Viktoriia Sharmanska is currently at Epic Games.

### v4

@@
-Efficient Online Inference of Vision Transformers by Training-Free Tokenization
+Visual-Word Tokenizer: Beyond Fixed Sets of Tokens in Vision Transformers
 Abstract
 The cost of deploying vision transformers increasingly represents a barrier to wider industrial adoption.
-Existing compression techniques require additional end-to-end fine-tuning or incur a significant drawback to runtime, making them ill-suited for online (real-time) inference, where a prediction is made on any new input as it comes in.
-We introduce the Visual Word Tokenizer (VWT), a training-free method for reducing energy costs while retaining performance and runtime.
-The VWT groups visual subwords (image patches) that are frequently used into visual words while infrequent ones remain intact.
+Existing compression techniques require additional end-to-end fine-tuning or incur a significant drawback to energy efficiency, making them ill-suited for online (real-time) inference, where a prediction is made on any new input as it comes in.
+We introduce the Visual-Word Tokenizer (VWT), a training-free method for reducing energy costs while retaining performance.
+The VWT groups visual subwords (image patches) that are frequently used into visual words, while infrequent ones remain intact.
 To do so, intra-image or inter-image statistics are leveraged to identify similar visual concepts for sequence compression.
-Experimentally, we demonstrate a reduction in wattage of up to 25% with only a 20% increase in runtime at most.
-Comparative approaches of 8-bit quantization and token merging achieve a lower or similar energy efficiency but exact a higher toll on runtime (up to 100% or more).
+Experimentally, we demonstrate a reduction in energy consumed of up to 47%.
+Comparative approaches of 8-bit quantization and token merging can lead to significantly increased energy costs (up to 500% or more).
 Our results indicate that VWTs are well-suited for efficient online inference with a marginal compromise on performance.
+The experimental code for our paper is also made publicly available [^1] https://github.com/wearepal/visual-word-tokenizer.
 {NoHyper}* Corresponding author: jg717@sussex.ac.uk.
-Figure 1: Visualization of patch matching by our visual word tokenizer (${\mathcal{T}}_{inter}^{100}$ model).
+Figure 1: Visualization of the inter-image approach by the Visual-Word Tokenizer (${\mathcal{T}}_{inter}^{100}$ model).
 Patches that are matched with one another are indicated by identical colors.
-Higher patch matching is exhibited by the background rather than foreground across datasets.
-Patch matching serves as a rudimentary form of image segmentation by grouping similar non-adjacent visual concepts.
+Higher patch matching is exhibited with the background rather than foreground object across the datasets.
+Patch matching serves as a rudimentary form of image segmentation by grouping similar non-adjacent visual concepts during inference.
 1 Introduction
@@
 However, due to the size and complexity of modern architectures, such models are costly to run in practice.
-To address this, various methods have been proposed to improve model efficiency such as Knowledge Distillation [24], Pruning [22, 41], and Quantization [11].
-However, many of these methods either require end-to-end fine-tuning to recover performance or significantly reduce runtime.
-In the field of Natural Language Processing (NLP), there is a growing trend towards improving efficiency via tokenization [18, 19, 9, 67, 43].
-Newer large language models (LLMs) [58, 14] exhibit a noticeably larger vocabulary than their earlier counterparts [12, 45], thereby producing shorter sequences across various distributions.
-For computer vision, increasing interest is placed on reducing the cost of deploying the vision transformer (ViT) [13].
+To address this, various methods have been proposed to improve model efficiency such as Knowledge Distillation [hinton2015distilling], Pruning [han2015learning, michel2019sixteen], and Quantization [dettmers2022gpt3].
+Many of these methods either require end-to-end fine-tuning to recover performance or significantly reduce energy efficiency.
+In the field of Natural Language Processing (NLP), there is a growing trend towards improving efficiency via tokenization [gee2022fast, gee2023multi, dagan2024getting, yamaguchi2024empirical, minixhofer2024zero].
+Newer large language models (LLMs) [team2024gemma, dubey2024llama] exhibit a noticeably larger vocabulary than their earlier counterparts [devlin2018bert, radford2019language], thereby producing shorter sequences across various distributions.
+For computer vision, increasing interest is placed on reducing the cost of deploying the vision transformer (ViT) [dosovitskiy2020image].
 As image encoders in larger vision-language systems, ViTs are used to process images as fixed sets of tokens.
-Similar to downsampling in convolutional neural networks, most research [27, 32, 48, 40, 5, 4, 26] has focused on merging and/or pruning tokens in the intermediate layers to reduce computational overhead.
-Given the analogous architecture of the transformer across image and text modalities, our work looks instead at the idea of tokenization for efficiency by splitting an image into variable sets of tokens.
-To introduce variability, we draw inspiration from subword tokenization algorithms [17, 51] used in NLP which follow the principle that common words should remain intact while infrequent ones are broken down into meaningful subword units.
+Similar to downsampling in convolutional neural networks, most research [kong2022spvit, liang2022not, rao2021dynamicvit, marin2021token, bolya2022token, bian2023multi, kim2024token] has focused on merging and/or pruning tokens in the intermediate layers to reduce computational overhead.
+Given the analogous architecture of the transformer across modalities, our work looks instead at the idea of tokenization for efficiency by splitting an image into variable sets of tokens.
+To introduce variability, we draw inspiration from subword tokenization algorithms [gage1994new, sennrich2016neural] used in NLP, which follow the principle that common words should remain intact while infrequent ones are broken down into meaningful subword units.
 Instead of a top-down approach – splitting words into subwords, our work for image data takes a bottom-up approach by grouping visual subwords (image patches) into visual words.
-We also twist the underlying principle: frequently used patches should be grouped as they are more likely to describe common features while infrequent ones remain intact as they might carry task-relevant information.
+We also twist the underlying principle: frequently used patches should be grouped as they are more likely to describe common features, while infrequent ones remain intact as they might carry task-relevant information.
 We propose two procedures to capture this principle.
-The first is an intra-image approach where patches with the lowest pixel variance within each image are grouped as they typically represent uniform areas (e.g.
-backgrounds).
+The first is an intra-image approach where patches with the lowest pixel variance within each image are grouped as they typically represent uniform areas (e.g., backgrounds).
 The second is an inter-image approach where basic features across multiple images such as colors or edges are discovered as visual words.
 Image patches are then grouped based on the similarity of these basic characteristics.
-Crucially, patches that have distinct characteristics (i.e.
-high dissimilarity with any visual word) remain intact and form separate visual subwords.
-The code for our paper is publicly available [^1] https://github.com/wearepal/visual-word-tokenizer.
+Crucially, patches that have distinct characteristics (i.e., high dissimilarity with any visual word) remain intact and form separate visual subwords.
 2 Related Work
@@
 Most works for improving the efficiency of ViTs have focused on reducing tokens in the intermediate layers by leveraging importance scores.
-In Liang et al.
-[32], Xu et al.
-[66], Kong et al.
-[27], Bian et al.
-[4], redundancy is addressed by fusing tokens.
-Both Rao et al.
-[48] and Tang et al.
-[55] opt to prune such tokens instead.
-Recent efforts [7, 8, 6, 26] attempt to combine the benefits of merging and pruning.
-In Tran et al.
-[60], an additional metric termed the energy score is used to better identify redundancy.
-Uniquely, Fayyaz et al.
-[16] use inverse transform sampling to select important tokens.
-Most relevant to our work are Marin et al.
-[40] and Bolya et al.
-[5].
-The former assigns tokens to centroids via clustering, while the latter progressively merges tokens layer-by-layer in a training-free manner [^2] Unlike Bolya et al.
-[5], we do not include Marin et al.
-[40] as one of our baselines due to a lack of code release..
-Figure 2: Overview of the Visual Word Tokenizer.
-An intra-image approach.
-In the forward pass, the pixel variance of the patches is computed with the top-k lowest values being masked (deploy).
-The grouped tokens are dropped after positional information is added.
-Note that ${\mathcal{V}}=1$ as only a single pixel variance feature is used.
-An inter-image approach.
-A Bag-of-Visual Words is formed by clustering patches in the pixel space (pre-process).
-In the forward pass, the minimum pairwise cosine distance between the patches and visual words is computed with values above the threshold being masked (deploy).
-The grouped tokens are averaged after positional information is added.
+In liang2022not, xu2022evo, kong2022spvit, bian2023multi, redundancy is addressed by fusing tokens.
+Both rao2021dynamicvit and tang2022patch opt to prune such tokens instead.
+Recent efforts [cao2023pumer, chen2023diffrate, bonnaerens2023learned, kim2024token] attempt to combine the benefits of merging and pruning.
+In tran2024accelerating, an additional metric termed the energy score is used to better identify redundancy.
+Uniquely, fayyaz2022adaptive use inverse transform sampling to select important tokens.
+Most relevant to our work are marin2021token and bolya2022token.
+The former assigns tokens to centroids via clustering, while the latter progressively merges tokens layer-by-layer in a training-free manner [^2] Unlike bolya2022token, we do not include marin2021token as one of our baselines due to a lack of code release..
+Figure 2: Overview of the Visual-Word Tokenizer (VWT).
+An intra-image approach (deploy only).
+During inference, the pixel variance of the patches is computed with the top-k lowest values being masked.
+The masked tokens are dropped after positional information is added.
+An inter-image approach (pre-process & deploy).
+First, a Bag-of-Visual Words is formed by clustering patches in the pixel space (pre-process).
+Then, during inference, the minimum pairwise cosine distance between the patches and visual words is computed with values above the threshold being masked (deploy).
+The unmasked tokens are averaged based on their grouping to the same visual word after positional information is added.
 Specialized Tokenizers.
@@
 Increasingly, the tokenizer’s vocabulary is specialized to reduce the input token length.
-In Gee et al.
-[18], domain adaptation of the tokenizer ensures fewer subword or character tokens are produced.
-Gee et al.
-[19] followed up by introducing n-grams for tokenization beyond the word-level boundary.
-In Dagan et al.
-[9], tokenizer specialization is also shown to accelerate the task of code generation with modern LLMs.
-Meanwhile, Yamaguchi et al.
-[67] analyzed the effectiveness of various vocabulary adaptation techniques for efficient cross-lingual inference.
-Recently, Minixhofer et al.
-[43] leveraged hypernetworks for zero-shot tokenizer transfer of newly domain-adapted vocabularies.
-Dataset Model $Base$ ${\mathcal{T}}_{intra}^{0.5}$ ${\mathcal{T}}_{inter}^{100}$ ${\mathcal{T}}_{inter}^{1000}$ ${\mathcal{T}}_{inter}^{10000}$
+In gee2022fast, domain adaptation of the tokenizer ensures fewer subword or character tokens are produced.
+gee2023multi followed up by introducing n-grams for tokenization beyond the word-level boundary.
+In dagan2024getting, tokenizer specialization is also shown to accelerate the task of code generation with modern LLMs.
+Meanwhile, yamaguchi2024empirical analyzed the effectiveness of various vocabulary adaptation techniques for efficient cross-lingual inference.
+Recently, minixhofer2024zero leveraged hypernetworks for zero-shot tokenizer transfer of newly domain-adapted vocabularies.
+Vector Quantization.
+The idea of discretizing continuous distributions has been explored in many works, most recently for image generation.
+yang2007evaluating leveraged clustering for learning a codebook that maps keypoint descriptors to discrete visual words.
+In wu2020visual and bao2021beit, discretization is applied as part of the modelling for ViTs.
+van2017neural learned discrete image representations by introducing the Vector Quantised-Variational Autoencoder (VQ-VAE) approach.
+esser2021taming and yu2021vector further improved upon the VQ-VAE by combining the expressiveness of transformers with an adversarial framework.
+Increasingly, vision-language models paired with codebooks conduct image synthesis autoregressively [ramesh2021zero, yu2022scaling, lu2022unified, lu2024unified, team2023gemini, team2024chameleon, sun2024autoregressive].
+Lastly, yang2022visual tackled disentangled representation learning and scene decomposition by tokenizing images into separate visual concepts.
+3 Visual-Word Tokenizer
+Dataset Model BaseBase ${\mathcal{T}}_{intra}^{0.5}$ ${\mathcal{T}}_{inter}^{100}$ ${\mathcal{T}}_{inter}^{1000}$ ${\mathcal{T}}_{inter}^{10000}$
 In-Domain ImageNet In-Domain ImageNet In-Domain ImageNet
@@
 ${\mathcal{T}}_{inter}^{{\mathcal{V}}}$ of varying pre-processing data and vocabulary sizes are shown.
-Unlike text tokenizers, domain specialization (i.e.
-In-Domain) does not result in greater compression.
+VWTs significantly reduce the token length of $Base$, particularly for larger image sizes (384 $\times$ 384 on COCO and NoCaps).
+Unlike text tokenizers, domain specialization (In-Domain) does not result in greater compression for the interinter-image approach.
 Smaller vocabularies produce shorter sequences whereas larger vocabularies result in patches being increasingly matched to separate visual words.
-Dataset Model $Base$ $Q_{8}$ $ToME$ ${\mathcal{T}}_{intra}^{0.5}$ ${\mathcal{T}}_{inter}^{100}$ ${\mathcal{T}}_{inter}^{1000}$ ${\mathcal{T}}_{inter}^{10000}$
-Waterbirds CLIP 123.99 73.89 91.64 110.93 102.65 107.93 117.36
-CelebA 126.53 74.47 93.07 114.37 96.31 102.12 115.46
-MetaShift 123.30 74.02 90.96 113.84 98.98 105.78 114.70
-OpenImages (Com.) 135.05 74.80 94.63 114.62 106.33 112.67 123.95
-OpenImages (Rare) 135.66 75.27 94.04 115.78 104.73 110.75 123.05
-COCO BLIP 191.74 81.68 147.23 169.77 144.71 156.25 171.84
-NoCaps 185.12 81.89 149.26 175.06 141.20 155.11 170.38
-(a) Wattage (watt) $\downarrow$
-Dataset Model $Base$ $Q_{8}$ $ToME$ ${\mathcal{T}}_{intra}^{0.5}$ ${\mathcal{T}}_{inter}^{100}$ ${\mathcal{T}}_{inter}^{1000}$ ${\mathcal{T}}_{inter}^{10000}$
-Waterbirds CLIP 8.48 88.55 17.04 5.84 9.88 9.73 9.91
-CelebA 8.33 89.41 17.25 5.86 9.45 9.54 9.99
-MetaShift 8.21 89.48 17.23 6.01 9.99 9.87 9.75
-OpenImages (Com.) 8.83 92.23 17.44 5.97 9.64 9.69 10.23
-OpenImages (Rare) 8.60 86.00 17.04 6.19 9.69 10.02 9.67
-COCO BLIP 12.45 57.11 15.06 8.10 8.73 9.10 10.31
-NoCaps 12.44 57.61 15.08 8.04 9.49 9.28 10.51
-(b) Runtime (millisecond) $\downarrow$
-Table 2: Wattage and runtime per sample.
-Compared to $Base$, VWTs reduce wattages by up to 25% with ${\mathcal{T}}_{inter}^{100}$.
-Although $Q_{8}$ and $ToME$ display lower wattages than VWTs, they induce significantly longer runtimes, especially the former.
-${\mathcal{T}}_{intra}^{0.5}$ is consistently faster than $Base$, while ${\mathcal{T}}_{inter}^{10000}$ may increase runtimes by up to 20% at most (except on COCO and NoCaps).
-Vector Quantization.
-The idea of discretizing continuous distributions has been explored in many works, most recently for image generation.
-Yang et al.
-[68] leveraged clustering for learning a codebook that maps keypoint descriptors to discrete visual words.
-In Wu et al.
-[64] and Bao et al.
-[3], discretization is applied as part of the modelling for ViTs.
-Van Den Oord et al.
-[61] learned discrete image representations by introducing the Vector Quantised-Variational Autoencoder (VQ-VAE) approach.
-Esser et al.
-[15] and Yu et al.
-[71] further improved upon the VQ-VAE by combining the expressiveness of transformers with an adversarial framework.
-Increasingly, vision-language models paired with codebooks conduct image synthesis autoregressively [47, 72, 38, 39, 57, 56, 54].
-Lastly, Yang et al.
-[69] tackled disentangled representation learning and scene decomposition by tokenizing images into separate visual concepts.
-3 Visual Word Tokenizer
+Dataset Model BaseBase $Q_{8}$ ToMeToMe ${\mathcal{T}}_{intra}^{0.5}$ ${\mathcal{T}}_{inter}^{100}$ ${\mathcal{T}}_{inter}^{1000}$ ${\mathcal{T}}_{inter}^{10000}$
+Waterbirds CLIP 1.05 6.54 1.56 0.65 1.01 1.05 1.16
+CelebA 1.05 6.66 1.61 0.67 0.91 0.97 1.15
+MetaShift 1.01 6.62 1.57 0.68 0.99 1.04 1.12
+OpenImages (Com.) 1.19 6.90 1.65 0.68 1.03 1.09 1.27
+OpenImages (Rare) 1.17 6.47 1.60 0.72 1.01 1.11 1.19
+COCO BLIP 2.39 4.66 2.22 1.38 1.26 1.42 1.77
+NoCaps 2.30 4.72 2.25 1.41 1.34 1.44 1.79
+Table 2: Energy (joule) consumed per sample.
+The power and runtime values that were used to compute the energy consumed are provided in Table 9.
+Compared to $Base$, VWTs reduce energy usage by up to 43% with ${\mathcal{T}}_{intra}^{0.5}$ and 47% with ${\mathcal{T}}_{inter}^{100}$.
+Both $Q_{8}$ and ToMeToMe significantly increase energy costs, particularly on datasets with smaller image sizes (224 $\times$ 224 on Waterbirds, CelebA, MetaShift, OpenImages).
+For the inter-image approach, efficiency improvements are highest with larger image sizes (384 $\times$ 384 on COCO and NoCaps) and decreases naturally with larger vocabularies due to smaller compression.
 In ViTs, tokenization is a process that splits an image into patches (tokens) which are then projected into a series of embeddings.
-The number of patches is typically fixed (e.g.
-$197$) based on the choice of architecture.
-We seek to split an image into variable length inputs instead (i.e.
-certain images will use $80$ tokens, while others a $100$).
-This variability is induced by conventional text tokenizers [17, 51, 65, 28] for model efficiency.
+The number of patches is typically fixed (e.g., $197$) based on the choice of architecture.
+We seek to split an image into variable length inputs instead (i.e., certain images will use $80$ tokens, while others a $100$).
+This variability is induced by conventional text tokenizers [gage1994new, sennrich2016neural, wu2016google, kudo2018sentencepiece] for model efficiency.
 We propose to achieve this via visual words that group patches (visual subwords) based on some commonality in the pixel space.
@@
 A simple yet effective grouping can be done using either a criterion that looks at statistics of only one image (an intra-image approach) or across many images (an inter-image approach).
-Figure 2 summarizes the Visual Word Tokenizer (VWT).
+Figure 2 summarizes the Visual-Word Tokenizer (VWT).
 Model Waterbirds CelebA MetaShift OpenImages
 Average $\uparrow$ Worst $\uparrow$ Average $\uparrow$ Worst $\uparrow$ Average $\uparrow$ Worst $\uparrow$ Common $\uparrow$ Rare $\uparrow$
-$Base$ 79.06 21.86 89.61 47.21 95.31 87.69 70.48 63.36
+BaseBase 79.06 21.86 89.61 47.21 95.31 87.69 70.48 63.36
 $Q_{8}$ 79.94 24.25 89.73 48.19 95.23 88.21 70.48 63.19
-$ToME$ 79.80 27.05 89.25 28.52 94.24 87.37 65.69 60.05
+ToMeToMe 79.80 27.05 89.25 28.52 94.24 87.37 65.69 60.05
 ${\mathcal{T}}_{intra}^{0.5}$ 75.56 31.00 89.27 50.93 92.94 86.15 65.52 59.73
@@
 BLEU@4 CIDEr CIDEr SPICE CIDEr SPICE CIDEr SPICE CIDEr SPICE
-$Base$ 34.00 107.35 103.57 14.60 98.87 14.11 92.96 13.30 98.34 14.02
+BaseBase 34.00 107.35 103.57 14.60 98.87 14.11 92.96 13.30 98.34 14.02
 $Q_{8}$ 33.85 106.87 103.86 14.53 100.07 14.18 92.72 13.49 99.12 14.10
-$ToME$ 30.02 94.40 97.35 14.42 90.74 13.39 84.22 12.69 90.37 13.40
+ToMeToMe 30.02 94.40 97.35 14.42 90.74 13.39 84.22 12.69 90.37 13.40
 ${\mathcal{T}}_{intra}^{0.5}$ 32.80 104.37 99.86 14.20 94.21 13.73 89.52 13.12 94.07 13.68
@@
 Table 3: Zero-shot (training-free) image classification (CLIP), subgroup robustness (CLIP) and captioning (BLIP).
-VWTs maintain average accuracy relative to $Base$ while improving robustness on Waterbirds.
-On COCO and NoCaps, VWTs retain a higher performance than $ToME$.
+$Q_{8}$ displays the lowest degradation in performance relative to $Base$, but suffers from a significant increase in energy used as shown in Table 2.
+The VWTs are competitive with $ToMe$, particularly for image captioning on COCO and NoCaps.
+Both $Q_{8}$ and the VWTs are shown to improve the worst-group accuracies on Waterbirds and CelebA.
+For the latter, the intraintra-image approach works best on MetaShift and OpenImages while the interinter-image approach is more suitable on the remaining datasets.
+The lower performance of the interinter-image approach on MetaShift and OpenImages may stem from the lack of uniform backgrounds (e.g., skies in Waterbirds) that increasingly shifts compression to the foreground object.
 3.1 An intra-image approach
@@
 We do not include the [CLS] token for dropping.
-This approach is inspired by Minderer et al.
-[42] which aimed to reduce the training cost of OWLv2, an open-vocabulary object detector.
+This approach is inspired by minderer2024scaling which aimed to reduce the training cost of OWLv2, an open-vocabulary object detector.
 Dropping patches with the lowest pixel variance removes padding areas and uniform backgrounds from the input mosaics of raw images, thereby increasing training efficiency.
@@
 Inspired by text tokenizers and codebooks, we propose a variable-length approach that statistically discovers visual words across many images.
-The tokenizer consists of two steps: Pre-Process and Deploy
-Figure 3: Visualization of patch dropping by our visual word tokenizer (${\mathcal{T}}_{intra}^{0.5}$ model).
+The tokenizer consists of two steps: Pre-Process and Deploy.
+Figure 3: Visualization of the intra-image approach by the Visual-Word Tokenizer (${\mathcal{T}}_{intra}^{0.5}$ model).
 Patches with the lowest pixel variance that are dropped are indicated in black.
-In most cases, the dropped patches correspond to the background object on average which is typically uninformative.
+In most cases, the dropped patches correspond to the uniform background which is uninformative.
+However, in a few cases, the patches of the foreground object are dropped (e.g., airplane in first image for COCO) which is undesirable.
+Figure 4: Visualization of the vocabulary of ${\mathcal{T}}_{inter}^{100}$.
+Patches from ImageNet-1K are matched to the closest visual word using the Euclidean distance.
+Visual words are shown to depict basic features such as colors or edges.
+Each visual word is an average representation of patches that belong to the matched cluster.
 Pre-Process.
 The Bag-of-Visual Words (BoVW) is a popular method for modeling images via discrete representations.
-In Yang et al.
-[68], k-means clustering is applied to keypoint descriptors from SIFT [37] to learn a fixed set of centroids.
+In yang2007evaluating, k-means clustering is applied to keypoint descriptors from SIFT [lowe2004distinctive] to learn a fixed set of centroids.
 These centroids represent the vocabulary to which multiple descriptors are mapped in a process termed Vector Quantization (VQ).
@@
 Second, in our early experimentation, we observed that patches in the embedding space have little similarity to one another.
-Such issues were also described by Bolya et al.
-[5], thus leading to their use of attention scores instead.
+Such issues were also described by bolya2022token, thus leading to their use of attention scores instead.
 Further justification for leveraging the pixel space is provided in Section 4.4.
-Given a dataset, we first patchify the images using the same patch size as the image encoder (e.g.
-16 for ViT-B/16).
+Given a dataset, we first patchify the images using the same patch size as the image encoder (e.g., 16 for ViT-B/16).
 We then cluster the patches via k-means to form the BoVW of a given vocabulary size, where each centroid represents a visual word.
-Patchification is done via basic tensor operations and not the pre-trained convolutional layer of the ViT to avoid projection into embeddings.
-We also execute this process in batches using the MiniBatchKMeans333from sklearn.cluster import MiniBatchKMeans algorithm due to memory constraints.
+Patchification is done via basic tensor operations [^3] We decompose the image into smaller tensors given the patch size while avoiding any linear transformations.
+and not the pre-trained convolutional layer of the ViT to avoid projection into embeddings.
+We also execute this process in batches using the MiniBatchKMeans444from sklearn.cluster import MiniBatchKMeans algorithm due to memory constraints.
 Note that MiniBatchKMeans uses the Euclidean distance by design.
 Since clustering is done in the pixel space, the BoVW may be reused by other ViTs with the same patch size regardless of model type.
+Figure 5: Visualization of long-form captioning on COCO.
+Longer captions are generated via a length penalty of 2.0 and a maximum length of 40.
+Interestingly, the smaller vocabulary of ${\mathcal{T}}_{inter}^{100}$ possesses higher descriptiveness and coherence than ${\mathcal{T}}_{intra}^{0.5}$, ${\mathcal{T}}_{inter}^{1000}$, or ${\mathcal{T}}_{inter}^{10000}$ despite its higher compression.
 Deploy.
@@
 We opt to append additional [PAD] tokens until the longest length within the batch is achieved.
-Similar to text transformers [62], the attention scores are set to negative infinity before the softmax to nullify the influence of padding.
+Similar to text transformers [vaswani2017attention], the attention scores are set to negative infinity before the softmax to nullify the influence of padding.
 We do not add positional information to the [PAD] tokens as extrapolating such information to non-uniform sequences will significantly worsen model efficiency.
@@
 The tokenizer transforms the input $x$ into tokens $t_{1},\ldots,t_{M}$, where $M\ll N$.
-In our experiments, we seek to analyze the effect of ${\mathcal{T}}$ on online inference efficiency.
+In our experiments, we seek to analyze the effect of $\mathcal{T}$} for online inference.
 We focus on the zero-shot setting by eschewing any form of end-to-end fine-tuning for $f$.
-Figure 4: Visualization of the vocabulary of ${\mathcal{T}}_{inter}^{100}$.
-Patches from ImageNet-1K are matched to the closest visual word using the Euclidean distance.
-Visual words are shown to depict basic features such as colors or edges.
-Each visual word is an average representation of patches that belong to the matched cluster.
+Model Waterbirds CelebA MetaShift OpenImages
+Average $\uparrow$ Worst $\uparrow$ Average $\uparrow$ Worst $\uparrow$ Average $\uparrow$ Worst $\uparrow$ Common $\uparrow$ Rare $\uparrow$
+${\mathcal{T}}_{intra}^{0.5}+Q_{8}$ 76.31 33.65 89.43 52.41 93.14 87.78 65.29 59.27
+${\mathcal{T}}_{inter}^{100}+Q_{8}$ 78.88 27.57 90.17 54.44 90.58 83.72 62.75 57.99
+${\mathcal{T}}_{inter}^{1000}+Q_{8}$ 80.01 25.60 91.05 50.93 93.67 86.67 66.15 60.48
+${\mathcal{T}}_{inter}^{10000}+Q_{8}$ 80.28 25.29 90.26 47.96 94.58 86.67 69.07 62.37
+(a) CLIP (image classification and subgroup robustness)
+Model COCO NoCaps
+Karpathy $\uparrow$ In-Domain $\uparrow$ Near-Domain $\uparrow$ Out-of-Domain $\uparrow$ Overall $\uparrow$
+BLEU@4 CIDEr CIDEr SPICE CIDEr SPICE CIDEr SPICE CIDEr SPICE
+${\mathcal{T}}_{intra}^{0.5}+Q_{8}$ 32.45 103.18 101.98 14.46 94.93 13.74 87.13 12.78 94.36 13.66
+${\mathcal{T}}_{inter}^{100}+Q_{8}$ 30.88 97.48 95.97 13.94 90.27 13.26 80.02 12.39 89.01 13.19
+${\mathcal{T}}_{inter}^{1000}+Q_{8}$ 31.72 100.23 99.69 14.02 94.35 13.49 85.69 12.57 93.36 13.39
+${\mathcal{T}}_{inter}^{10000}+Q_{8}$ 32.98 104.15 101.77 14.34 98.75 14.02 91.55 13.21 97.72 13.91
+(b) BLIP (image captioning)
+Table 4: Zero-shot (training-free) image classification (CLIP), subgroup robustness (CLIP) and captioning (BLIP).
+The VWTs are applied jointly with 8-bit quantization.
+Performance is shown to be similar to those with VWTs only, thereby displaying the mutual compatibility of VWTs with other compression techniques.
 4.1 Datasets and Settings
 We conduct our analysis through the lens of (i) classification performance of visual recognition, (ii) subgroup robustness, and (iii) generative performance of visual captioning.
-For (i) and (ii), we utilize three publicly available datasets (Waterbirds [63], CelebA [36], MetaShift [31]) that are typical benchmarks in robustness and fairness research [50, 35, 70].
-To perform (zero-shot) classification, we compute the cosine similarity between an image embedding and the following encoded text labels4 for Waterbirds, CelebA, and MetaShift, respectively: {’landbird’, ’waterbird’}, {’non-blond’, ’blond’}, {’dog’, ’cat’}.
+For (i) and (ii), we utilize three publicly available datasets (Waterbirds [wah2011caltech], CelebA [liu2015deep], MetaShift [liang2022metashift]) that are typical benchmarks in robustness and fairness research [sagawa2019distributionally, liu2021just, yang2023change].
+To perform (zero-shot) classification, we compute the cosine similarity between an image embedding and the following encoded text labels5 for Waterbirds, CelebA, and MetaShift, respectively: {’landbird’, ’waterbird’}, {’non-blond’, ’blond’}, {’dog’, ’cat’}.
 Further details on the defined subgroups are provided in Appendix 6.1.
-For (i), we also conduct a large-scale evaluation on the OpenImages v6 dataset [29].
-Following Huang et al.
-[25], the test split is divided into common and rare subsets that consist of 57 224 and 21 991 images, respectively.
+For (i), we also conduct a large-scale evaluation on the OpenImages v6 dataset [kuznetsova2020open].
+Following huang2023open, the test split is divided into common and rare subsets that consist of 57 224 and 21 991 images, respectively.
 The former has 214 classes, while the latter has 200.
 To perform (zero-shot) classification, we compute the cosine similarity between an image embedding and the encoded text label [^5] The prefix ”a photo of a ” is also added to encode each text label..
-For (iii), we utilize the Karpathy test split of COCO dataset [33] and a validation set of NoCaps dataset [1] following the setting in previous work [30].
-Finally, to study inference efficiency, we utilize all datasets for the visual tasks (i)-(iii) described above.
-We compute the efficiency per sample and average across all samples.
-For the wattage calculation, we call ”pynvml.nvmlDeviceGetPowerUsage(handle)/1000” as done by Weights and Biases, which leverages the NVIDIA Management Library for monitoring and managing the NVIDIA GPUs.
+For (iii), we utilize the Karpathy test split of COCO dataset [lin2014microsoft] and a validation set of NoCaps dataset [agrawal2019nocaps] following the settings in previous work [li2022blip].
+Finally, to study inference efficiency, we utilize all the datasets for the visual tasks (i)-(iii) described above in computing the energy (joule) consumed per sample.
+For calculating the power variable of energy, we call ”pynvml.nvmlDeviceGetPowerUsage(handle)/1000” as done by Weights and Biases, which leverages the NVIDIA Management Library for monitoring and managing the NVIDIA GPUs.
 The experiments are also conducted in an isolated environment, thus ensuring the efficiency measurements are as accurate as possible.
-Note that we do not include the Pre-Process step of the inter-image approach into our efficiency calculations as this process is only done once and may be reused multiple times.
+Note that we do not include the Pre-Process step of the inter-image approach into our efficiency calculations, as this process is only done once and may be reused multiple times.
 4.2 Implementation Details
-For image classification, we load the pre-trained CLIP [46] model from HuggingFace [^5] https://huggingface.co/openai/clip-vit-base-patch16.
+For image classification, we load the pre-trained CLIP [radford2021learning] model from HuggingFace [^6] https://huggingface.co/openai/clip-vit-base-patch16.
 An image size of 224 $\times$ 224 is used with bilinear interpolation for CLIP.
-For image captioning, we load the pre-trained BLIP [30] model from HuggingFace [^6] https://huggingface.co/Salesforce/blip-image-captioning-base.
+For image captioning, we load the pre-trained BLIP [li2022blip] model from HuggingFace [^7] https://huggingface.co/Salesforce/blip-image-captioning-base.
 To perform zero-shot captioning, we use a beam size of 3 along with maximum and minimum lengths of 20 and 5, respectively.
@@
 Both CLIP and BLIP utilize the ViT-B/16 image encoder unless stated otherwise.
-Aside from the pre-trained model which we denote as $Base$, we also consider 8-bit quantization [11] and token merging [5] as additional baselines.
+Aside from the pre-trained model which we denote as $Base$, we also consider 8-bit quantization [dettmers2022gpt3] and token merging [bolya2022token] as additional baselines.
 We denote the former as $Q_{8}$ and the latter as $ToMe$.
-Following Bolya et al.
-[5], we utilize a reduction per layer for $ToME$ of 13 with CLIP and 23 with BLIP due to their respective input image sizes.
-For the VWTs, we set the top-k of the intra-image approach to 50% of the total number of patches which we denote as ${\mathcal{T}}_{intra}^{0.5}$.
+Following bolya2022token, we utilize a reduction per layer for $ToMe$ of 13 with CLIP and 23 with BLIP due to their respective input image sizes.
+For the VWTs, we set the top-k of the intra-image approach to 50% of the total number of patches which we denote as ${\mathcal{T}}_{intra}^{0.5}$ as we found it to work best.
+Additional dropping ratios are explored in Table 10 of Appendix 7.1.
 For the inter-image approach, we set the threshold to 0.1 unless stated otherwise and denote it as ${\mathcal{T}}_{inter}^{{\mathcal{V}}}$, where $\mathcal{V}$} is the size of the vocabulary.
@@
 Since our focus is on the online setting (real-time), we set the batch size to 1 unless stated otherwise.
-Figure 5: Visualization of long-form captioning on COCO.
-Longer captions are generated via a length penalty of 2.0 and a maximum length of 40.
-Interestingly, the smaller vocabulary of ${\mathcal{T}}_{inter}^{100}$ possesses higher descriptiveness and coherence than ${\mathcal{T}}_{intra}^{0.5}$, ${\mathcal{T}}_{inter}^{1000}$, or ${\mathcal{T}}_{inter}^{10000}$ despite its lesser compression.
-4.3 VWTs and Inference Efficiency
-First we analyze the effects of VWTs on token length.
-We seek to understand how the choice of pre-processing data and vocabulary size affects the degree of compression.
-Table 1 shows the token length per sample (including [CLS]) on different datasets.
-Unlike ${\mathcal{T}}_{intra}^{0.5}$, the sequence lengths induced by ${\mathcal{T}}_{inter}^{{\mathcal{V}}}$ are not equal.
-First, we compare ${\mathcal{T}}_{inter}^{{\mathcal{V}}}$ pre-processed on the in-domain dataset and ImageNet-1K [10].
-The in-domain dataset is represented by the training split if available.
-On text data, in-domain tokenizers [18, 19, 9, 67, 43] have been shown to produce shorter sequences by specializing the vocabulary on the given distribution.
-Interestingly, we observe no such effect with image data as seen by the similar lengths between In-Domain and ImageNet-1K.
-Only on CelebA, do we see a slightly greater reduction with ${\mathcal{T}}_{inter}^{1000}$ and ${\mathcal{T}}_{inter}^{10000}$ pre-processed on ImageNet-1K.
-Second, unlike text tokenizers, decreasing compression is seen as vocabulary size increases.
-With text, larger vocabularies ensure that more tokens are kept as words rather than subwords.
-We posit that an increasing number of patches are matched to separate visual words, thus lowering the overall compression.
-Having analyzed the effects on token length, we turn to the practical metrics of wattage and runtime.
-For online inference, savings in wattage should not incur a significant cost to runtime.
-In Table 2, we compare the efficiency of VWTs to $Base$, $Q_{8}$, and $ToME$.
-Note that we compute both metrics using the image encoder of CLIP or BLIP only.
-First, we find the wattage of VWTs to be lower than $Base$ across the datasets.
-On COCO, this reduction is up to 25% with ${\mathcal{T}}_{inter}^{100}$.
-Naturally, efficiency decreases as vocabulary size increases due to smaller compression.
-Although $Q_{8}$ and $ToME$ result in a lower wattage than VWTs, we note the drawbacks in runtime.
-In particular, $Q_{8}$ results in a significantly longer runtime for $Base$ as tensors need to be repeatedly quantized and dequantized.
-Meanwhile, $ToME$ can increase runtime by up to 100% on Waterbirds, CelebA, MetaShift, and OpenImages.
-We observe increases of up to 20% at most with ${\mathcal{T}}_{inter}^{10000}$ on the aforementioned datasets.
-${\mathcal{T}}_{intra}^{0.5}$ displays lower runtimes than $Base$ across the datasets while ${\mathcal{T}}_{inter}^{100}$, ${\mathcal{T}}_{inter}^{1000}$, and ${\mathcal{T}}_{inter}^{10000}$ do so on COCO and NoCaps.
-We have shown how VWTs are more suitable for online inference by striking a good balance between savings in wattage and costs to runtime.
 Dataset Model ${\mathcal{T}}_{inter}^{100}$ ${\mathcal{T}}_{inter}^{1000}$ ${\mathcal{T}}_{inter}^{10000}$
@@
 NoCaps - 394 - 396 - 431
-Table 4: Token length per sample (including [CLS]).
+Table 5: Ablation of token length per sample (including [CLS]).
 ${\mathcal{T}}_{inter}^{{\mathcal{V}}}$ of varying pre-processing data and vocabulary sizes are shown.
-Visual words are formed from patches after the pre-trained convolution layer of CLIP or BLIP.
-Unlike Table 1, poor compression is seen due to dissimilarity between patches and new visual words.
-4.4 VWTs and Visual Performance
-Another important factor in compression is the effect on model performance.
-In Table 3, we tabulate the performance of image classification and captioning using CLIP and BLIP, respectively.
-For visual recognition and subgroup robustness, we analyze performance from an average and worst-group perspective as done by Sagawa et al.
-[50] and Romiti et al.
-[49].
-First, we find the degradation in average accuracy to be small for the VWTs.
-The largest drop of up to 2% only is shown by ${\mathcal{T}}_{intra}^{0.5}$ on MetaShift.
-Likewise, both $Q_{8}$ and $ToME$ retain a high average accuracy across the datasets.
-Second, we observe possible improvements in subgroup robustness with VWTs.
-On Waterbirds and CelebA, the worst-group accuracy (WGA) with ${\mathcal{T}}_{intra}^{0.5}$ increases by up to 29% and 8%, respectively.
-Only on MetaShift do we observe a lower WGA than $Q_{8}$ or $ToME$ relative to $Base$.
-Like Gee et al.
-[20], we find compression to not always be harmful to subgroup robustness by improving the WGA at a negligible cost to overall performance.
-To further validate performance, we conduct additional evaluations on the large-scale OpenImages v6 dataset for zero-shot classification.
-We report mean Average Precision (mAP) as done by Huang et al.
-[25].
-Likewise, from Tables 2 and 3, we find the VWTs to be more suitable for online inference than $Q_{8}$ or $ToME$ in the large-scale setting while retaining comparable performance to $Base$.
-For image captioning with BLIP, we evaluate our models following the setting in Li et al.
-[30] by using the BLEU, CIDEr, and SPICE metrics w.r.t.
-the ground truth captions.
-On COCO, the VWTs display a higher performance than $ToME$ and are competitive with $Q_{8}$.
-On NoCaps, we see the largest degradation on the out-of-domain samples with ${\mathcal{T}}_{inter}^{100}$.
-However, overall performance of ${\mathcal{T}}_{intra}^{0.5}$, ${\mathcal{T}}_{inter}^{1000}$, and ${\mathcal{T}}_{inter}^{10000}$ are still higher than $ToME$.
-Only $Q_{8}$ displays a slight improvement over $Base$.
-Like $ToME$, the VWTs are shown to not require additional end-to-end fine-tuning for performance retention.
-The generated captions by BLIP [30] can be used as priors for further training of other models.
-As such, longer descriptive captions may be more desirable than the typical short captions associated with Internet data.
-In Figure 5, we visualize the long-form captions on COCO.
-To enable longer generations, we set the length penalty to 2.0 and double the maximum length to 40.
-All other generation settings are kept the same.
-With longer captions, the generation may degenerate into unnecessary repetitions on certain samples.
-Interestingly, descriptiveness and coherence improves more with ${\mathcal{T}}_{inter}^{100}$ than ${\mathcal{T}}_{intra}^{0.5}$, ${\mathcal{T}}_{inter}^{1000}$, or ${\mathcal{T}}_{inter}^{10000}$ inspite of its higher sequence compression as seen on COCO in Table 1.
+Visual words for the inter-image approach are formed in the embedding space from patches after the pre-trained convolution layer of CLIP or BLIP.
+Unlike Table 1, poor compression is seen due to dissimilarity between the patches and new visual words during inference.
+4.3 Experimental Results
+VWTs and Inference Efficiency.
+We begin by analyzing the effects of VWTs on token length to understand how the choice of pre-processing data and vocabulary size affects the degree of compression.
+Table 1 shows the token length per sample (including [CLS]) on different datasets.
+First, we observe a significant reduction in token length by the VWTs relative to $Base$.
+This reduction is most apparent for datasets with larger image sizes such as COCO and NoCaps.
+On the former, the token length is reduced by up to 54% with ${\mathcal{T}}_{inter}^{100}$.
+Second, the token lengths induced by ${\mathcal{T}}_{inter}^{{\mathcal{V}}}$ are not equal unlike ${\mathcal{T}}_{intra}^{0.5}$.
+We compare ${\mathcal{T}}_{inter}^{{\mathcal{V}}}$ pre-processed on the in-domain dataset and ImageNet-1K [deng2009imagenet].
+The in-domain dataset is represented by the training split if available.
+On text data, in-domain tokenizers [gee2022fast, gee2023multi, dagan2024getting, yamaguchi2024empirical, minixhofer2024zero] have been shown to produce shorter sequences by specializing the vocabulary on the given distribution.
+Interestingly, we observe no such effect with image data as seen by the similar lengths between In-Domain and ImageNet-1K.
+Only on CelebA, do we see a slightly greater reduction with ${\mathcal{T}}_{inter}^{1000}$ and ${\mathcal{T}}_{inter}^{10000}$ pre-processed on ImageNet-1K.
+Third, unlike text tokenizers, decreasing compression is seen as vocabulary size increases.
+With text, larger vocabularies ensure that more tokens are kept as words rather than subwords.
+We posit that an increasing number of patches are matched to separate visual words, thus lowering the overall compression.
+Note that henceforth, we use ImageNet-1K as our pre-processing data for the inter-image approach.
 Dataset Model ${\mathcal{T}}_{inter}^{100}$ ${\mathcal{T}}_{inter}^{1000}$ ${\mathcal{T}}_{inter}^{10000}$
@@
 NoCaps
-(a) Length
+(a) Token Length
 Model Waterbirds CelebA MetaShift Overall $\uparrow$
@@
 (c) BLIP (image captioning)
-Table 5: Token length (including [CLS]) and performance of the inter-image approach with random merging.
+Table 6: Ablation of token length (including [CLS]) and performance of the inter-image approach with random merging.
 The pairwise cosine distance is initialized by sampling from a uniform distribution of $[0,2]$.
 Average and worst-group accuracies degrade noticeably with ${\mathcal{T}}_{inter}^{100}$ except for CelebA.
-Performance does not change much for ${\mathcal{T}}_{inter}^{1000}$ and ${\mathcal{T}}_{inter}^{10000}$ from $Base$ since barely any compression occurs.
-4.5 Visualization of the Visual Words
-To better understand the VWT, we visualize the patch matching of ${\mathcal{T}}_{intra}^{0.5}$ and ${\mathcal{T}}_{inter}^{100}$ in Figures 3 and 1, respectively.
-For the latter, we highlight in identical colors patches that are matched with one another.
-With ${\mathcal{T}}_{intra}^{0.5}$, the patches with the lowest pixel variance typically correspond to uninformative backgrounds.
-In most cases, the dropping of such patches continues to preserve the foreground object.
+Performance does not change significantly for ${\mathcal{T}}_{inter}^{1000}$ and ${\mathcal{T}}_{inter}^{10000}$ from $Base$ as negligible compression occurs.
+Having analyzed the effects on token length, we turn to the practical metrics of energy.
+In Table 2, we compare the efficiency of VWTs to $Base$, $Q_{8}$, and $ToMe$.
+Note that we compute both metrics using the image encoder of CLIP or BLIP only.
+First, we find the energy consumed by VWTs to be generally lower than $Base$ across the datasets.
+On OpenImages (Common) and COCO, this reduction is up to 43% with ${\mathcal{T}}_{intra}^{0.5}$ and 47% with ${\mathcal{T}}_{inter}^{100}$, respectively.
+Interestingly, the inter-image approach appears to be most effective for datasets with larger image sizes (384 $\times$ 384) as shown with COCO and NoCaps rather than the remaining datasets with smaller image sizes (224 $\times$ 224).
+This efficiency also decreases as the vocabulary size increases due to smaller compression.
+For $Q_{8}$ and $ToMe$, we observe significant drawbacks in energy efficiency relative to $Base$.
+In the case of the former, the increase in energy consumed can reach as high as 500% or more on Waterbirds, CelebA, and MetaShift.
+This can be explained by the significantly longer runtime of $Q_{8}$ as tensors need to be repeatedly quantized and dequantized.
+Meanwhile, $ToMe$ can increase energy consumption by up to 50% or more.
+We have shown how VWTs are more suitable for online inference by improving energy efficiency, particularly for datasets with larger image sizes.
+As mentioned previously, we do not consider the Pre-Process step of the inter-image approach in our efficiency calculations as it is only executed once and not repeated during inference.
+This is similar to how model efficiency for large language models is calculated by excluding the construction of the tokenizer’s vocabulary via Byte-Pair Encoding.
+For reference, this step on the CPU takes approximately 30 minutes to an hour, depending on the vocabulary size [^8] Note that this pre-processing is done on ImageNet-1K.
+Naturally, dataset size also affects the total duration..
+VWTs and Visual Performance.
+Another important factor in compression is the effect on model performance.
+In Table 3, we tabulate the performance of image classification and captioning using CLIP and BLIP, respectively.
+For visual recognition and subgroup robustness, we analyze performance from an average and worst-group perspective as done by sagawa2019distributionally and romiti2022realpatch.
+To further validate performance, we report the mean Average Precision (mAP) on the large-scale OpenImages v6 dataset [huang2023open].
+For image captioning with BLIP, we evaluate our models following the setting in li2022blip by using the BLEU, CIDEr, and SPICE metrics w.r.t.
+the ground truth captions.
+First, we find the degradation in performance relative to $Base$ to be smallest for $Q_{8}$ across the datasets.
+However, this comes at the cost of a significant rise in energy consumed as shown in Table 2.
+We also observe that VWTs are competitive with $ToMe$, particularly on the image captioning tasks of COCO and NoCaps.
+Second, both $Q_{8}$ and the VWTs are shown to improve the subgroup robustness on Waterbirds and CelebA.
+The worst-group accuracy (WGA) with ${\mathcal{T}}_{intra}^{0.5}$ increases by up to 29% and 8%, respectively.
+Only on MetaShift do we observe a lower WGA than $Q_{8}$ or $ToMe$.
+Like gee2023compressed, we find compression to not always be harmful to subgroup robustness by improving the WGA at a small or negligible cost to overall performance.
+Third, the intraintra-image approach is seen to work best on MetaShift and OpenImages while the interinter-image approach is more suitable on the remaining datasets.
+The average accuracy with ${\mathcal{T}}_{inter}^{100}$ drops by up to 5% and 7% on MetaShift and OpenImages, respectively.
+We posit that the lower performance may stem from the lack of uniform backgrounds (e.g., skies in Waterbirds) that causes compression to increasingly shift to the foreground object instead as shown in Figure 1.
+Model Waterbirds CelebA MetaShift
+Length Average $\uparrow$ Worst $\uparrow$ Length Average $\uparrow$ Worst $\uparrow$ Length Average $\uparrow$ Worst $\uparrow$
+0.2 97 76.71 24.09 66 89.44 56.48 84 86.16 77.46
+0.3 79 74.85 21.13 55 88.85 58.70 69 81.16 67.77
+0.4 66 73.35 16.30 50 88.31 57.04 60 77.73 60.65
+0.5 58 72.81 13.14 48 88.21 55.74 55 74.75 56.08
+Table 7: Ablation of token length (including [CLS]) and performance using ${\mathcal{T}}_{inter}^{100}$ with varying similarity thresholds.
+Higher thresholds for the inter-image approach can reduce sequences to 48 tokens on CelebA.
+Greater compression results in higher degradation in performance, particularly in subgroup robustness.
+Visualizing the VWTs.
+To better understand the VWTs, we visualize the patch dropping of ${\mathcal{T}}_{intra}^{0.5}$ and patch matching of ${\mathcal{T}}_{inter}^{100}$ in Figures 3 and 1, respectively.
+For the latter, we highlight in identical colors the patches that are matched with one another.
+With ${\mathcal{T}}_{intra}^{0.5}$, patches with the lowest pixel variance typically correspond to uninformative backgrounds.
+In most cases, dropping such patches continues to preserve the foreground object.
 With ${\mathcal{T}}_{inter}^{100}$, we find that patches representing the background are more frequently grouped than those of the foreground object.
-On Waterbirds, the merging of background patches may explain the improved robustness in Table 3 by mitigating the spurious correlation with the background.
-On CelebA, ${\mathcal{T}}_{inter}^{100}$ tends to avoid matching the eyes or mouths of the individuals.
-We also observe that patch matching is capable of grouping similar but non-adjacent visual concepts.
-In certain examples of MetaShift and NoCaps, multiple cats and foods are seemingly matched together, respectively.
+On Waterbirds, the merging of background patches may explain the improved robustness in Table 3 by mitigating spurious correlations with the background.
+On CelebA, ${\mathcal{T}}_{inter}^{100}$ tends to avoid matching the eyes or mouths.
+We also observe that ${\mathcal{T}}_{inter}^{100}$ may group similar but non-adjacent visual concepts.
+In certain examples of MetaShift and NoCaps, multiple cats and foods are matched together, respectively.
 Our analysis shows that patch matching serves as a rudimentary form of image segmentation by identifying similar visual concepts for compression.
 In Figure 4, we visualize the vocabulary of ${\mathcal{T}}_{inter}^{100}$ to analyze the formation of the visual words.
-We show patches from ImageNet-1K (i.e.
-pre-processing data) that are matched to each visual word using the Euclidean distance.
+We show patches from ImageNet-1K (pre-processing data) that are matched to each visual word using the Euclidean distance.
 Since visual words are centroids, patches that are matched to the same visual word belong to the same cluster.
-We observe that visual words depict basic features such as colors or edges.
+We find that visual words depict basic features such as colors or edges.
 These features are also formed as an average representation of the patches that belong to each cluster.
 By representing basic features, visual words serve as effective anchor points to which patches can be matched to.
-Dataset Thresh Length Average $\uparrow$ Worst $\uparrow$
-Waterbirds 0.2 97 76.71 24.09
-0.3 79 74.85 21.13
-0.4 66 73.35 16.30
-0.5 58 72.81 13.14
-CelebA 0.2 66 89.44 56.48
-0.3 55 88.85 58.70
-0.4 50 88.31 57.04
-0.5 48 88.21 55.74
-MetaShift 0.2 84 86.16 77.46
-0.3 69 81.16 67.77
-0.4 60 77.73 60.65
-0.5 55 74.75 56.08
-Table 6: Token length (including [CLS]) and performance using ${\mathcal{T}}_{inter}^{100}$ with varying similarity thresholds.
-Higher thresholds can reduce sequences to 48 tokens on CelebA.
-Greater compression results in higher degradation in performance, particularly in subgroup robustness.
-4.6 Ineffectiveness of the Embedding Space
-In Table 4, we analyze the sequence compression using visual words formed in the embedding space.
+The generated captions by BLIP [li2022blip] can be used as priors for further training of other models.
+As such, longer descriptive captions may be more desirable than the typical short captions associated with Internet data.
+In Figure 5, we visualize the long-form captions on COCO.
+To enable longer generations, we set the length penalty to 2.0 and double the maximum length to 40.
+All other generation settings are kept the same.
+With longer captions, the generation may degenerate into unnecessary repetitions on certain samples.
+Interestingly, descriptiveness and coherence improves more with ${\mathcal{T}}_{inter}^{100}$ than ${\mathcal{T}}_{intra}^{0.5}$, ${\mathcal{T}}_{inter}^{1000}$, or ${\mathcal{T}}_{inter}^{10000}$ inspite of its higher sequence compression as seen on COCO in Table 1.
+4.4 Ablation Studies
+VWTs with Quantization.
+We designed the VWT to be a training-free approach to efficient online inference.
+As such, we exclude any form of fine-tuning [lin2024not, hao2024meft, simoulin2024memory] (e.g., knowledge distillation [tian2019contrastive, han2024amd, son2021densely]) for performance recovery.
+However, should it be desired, our method may be used alongside network compression [tian2019contrastive, han2024amd, son2021densely] (i.e., minimizing the Kullback–Leibler divergence between the softened outputs of the larger teacher and smaller student), memory-efficient fine-tuning [hao2024meft, simoulin2024memory], and other compression techniques (e.g., quantization) as the VWT only influences the initial input sequence to the vision transformer.
+In Table 4, we provide additional scores when VWTs are combined with 8-bit quantization.
+We find the performance to be similar to those in Table 3, thereby showing that VWTs are highly compatible with other compression approaches as mentioned above.
+Figure 6: Visualization of image captions on COCO by the inter-image approach.
+The generated captions are shown to deviate more when increasing the similarity threshold than when reducing the vocabulary size.
+With random matching, the model begins to completely misunderstand the image.
+Embeddings as Visual Words.
+In Table 5, we analyze compression by the inter-image approach using visual words formed in the embedding space.
 Instead of initializing the vocabulary by clustering patches in the pixel space, we do so with patches after the pre-trained convolution layer of CLIP or BLIP.
@@
 Compared to Table 1, we observe a notable reduction in the compression irrespective of pre-processing data and vocabulary size.
-In Tang et al.
-[55], embeddings are shown to become progressively more similar to one another due to self-attention.
-As such, it is unsurprising that matching patches and visual words in the initial embedding space is found to be ineffective.
-4.7 Random Merging of Tokens
-In Table 5, we study the effects of randomly merging the tokens.
+In tang2022patch, embeddings are shown to become progressively more similar to one another due to self-attention.
+As such, it is unsurprising that matching patches and visual words in the initial embedding space (i.e., before any self-attention is applied) is found to be ineffective.
+Raising the Similarity Threshold.
+We have shown how the inter-image approach can improve the efficiency of online inference.
+To better understand its limitations, we ablate the similarity threshold of ${\mathcal{T}}_{inter}^{100}$ by setting it to values of $\{0.2,0.3,0.4,0.5\}$ in Table 7.
+We seek to determine if exploiting higher thresholds for increased compression is a viable strategy.
+Naturally, we observe a reduction in performance as increasingly dissimilar patches are merged.
+For Waterbirds and MetaShift, the WGA degrades more significantly than the average, especially on the former.
+Interestingly, average accuracy remains relatively unchanged while WGA improves significantly on CelebA irrespective of similarity threshold.
+We posit that at higher thresholds, the inter-image approach begins to merge core features that represent the foreground object, resulting in the reduced performance of Waterbirds and MetaShift during inference.
+Random Merging of Tokens.
+In Table 6, we study the effects of randomly merging tokens to contrast the Bag-of-Visual-Words criterion for the inter-image approach.
 We initialize the pairwise cosine distance by sampling from a uniform distribution of $[0,2]$.
@@
 For ${\mathcal{T}}_{inter}^{1000}$ and ${\mathcal{T}}_{inter}^{10000}$, performance does not shift much from $Base$ as barely any compression occurs.
-With random merging, the captions are shown to deviate completely from those of $Base$ in Figure 6, thus further demonstrating that visual words are an effective criterion for grouping patches as shown in Figure 1.
-4.8 Ablation of the Similarity Threshold
-We have shown how the inter-image approach can improve the efficiency of online inference.
-To better understand their limitations, we ablate the threshold of ${\mathcal{T}}_{inter}^{100}$ by setting it to values of $\{0.2,0.3,0.4,0.5\}$ in Table 6.
-We seek to determine if exploiting higher thresholds for increased compression is a viable strategy.
-Naturally, we observe a reduction in performance as increasingly dissimilar patches are merged.
-For Waterbirds and MetaShift, the WGA degrades more significantly than the average, especially with the former.
-Interestingly, average accuracy remains relatively unchanged while WGA improves significantly on CelebA irrespective of similarity threshold.
-We posit that at higher thresholds, the merging of core features represented by the foreground object results in the reduced performance of Waterbirds and MetaShift.
-Figure 6: Visualization of image captions on COCO by the inter-image approach.
-The generated captions are shown to deviate more when increasing the similarity threshold than when reducing the vocabulary size.
-With random matching, the model begins to completely misunderstand the image.
-4.9 When to use the intra-image or inter-image approach?
-Concerning the choice between the intra-image or inter-image approach, we posit that when global information is required, dropping tokens may be more advantageous by removing unnecessary noise from the visual input.
-On the other hand, when the task necessitates local information (e.g.
-long-form captioning), merging tokens may better preserve the visual concepts.
+With random matching, the captions are shown to deviate completely from those of $Base$ in Figure 6, thus further demonstrating that visual words are an effective criterion for grouping patches as shown in Figure 1.
+5 Conclusion
+In this work, we set out to define a training-free tokenization for ViTs that lowers the energy consumed while balancing costs to performance.
+In online scenarios, we have shown empirically that our intra-image and inter-image approaches are more suitable than 8-bit quantization and token merging for image classification and captioning.
+Visually, the criterion of the intra-image approach typically corresponds to the background, while that of the inter-image approach groups analogous visual concepts based on visual words that represent basic features.
+Concerning the choice between the intra-image or inter-image approach, when global information (e.g., image classification) is required, dropping tokens may be more advantageous by removing unnecessary noise from the visual input.
+On the other hand, when the task necessitates local information (e.g., long-form captioning), merging tokens may better preserve the visual concepts.
+This is also a limitation of our training-free approach that could restrict its viability in sensitive domains such as medical imaging or autonomous driving.
 For example, the image on the first row of Figure 7 of Appendix 7.1 shows that ${\mathcal{T}}_{intra}^{0.5}$ removes the mountainous background, thus leading to the absence of ”mountain” in the long-form caption of Figure 5.
-Potentially, combining both approaches may maximize compression while preserving information by merging the mountains and dropping the bushes.
-5 Conclusion
-In this work, we set out to define a training-free tokenization for ViTs that lowers wattage while balancing costs to runtime and performance.
-In online scenarios, we have shown empirically that our intra-image and inter-image approaches are stronger than 8-bit quantization and token merging for image classification and captioning.
-Analysis on large-scale classification further validates the viability of our method while long-form captioning shows its potential for improving descriptiveness and coherence.
-Qualitatively, we observe how the criterion of the intra-image approach typically corresponds to the background while that of the inter-image approach groups analogous visual concepts based on visual words that represent basic features.
-As a future work, additional research could explore combining both the intra-image and inter-image approaches, potentially improving further the performance and efficiency of the VWT.
+As a future work, combining both approaches may maximize compression while safeguarding against the loss of critical information [^9] For the current approach, a possible safeguard is to carefully tune the hyperparameters of the VWT (i.e., top-k or similarity threshold) to one’s requirements such that foreground objects are not unintentionally affected.
+such as by merging the mountains and dropping the bushes.
 Acknowledgements
@@
 Viktoriia Sharmanska is currently at Epic Games.
+6 Further Details
+6.1 Datasets
+Here, we detail the task of each dataset for subgroup robustness. In Table 8, we also tabulate the labels and attributes that define each subgroup along with their sample sizes.
+Waterbirds.
+Given an image of a bird, the task is to predict whether it is a waterbird or landbird [wah2011caltech].
+Following sagawa2019distributionally, the attribute is the background that the bird is on.
+We use the same dataset splits as liu2021just.
+CelebA.
+Given an image of a person, the task is to predict whether the hair color is blond or not [liu2015deep].
+Following sagawa2019distributionally, the attribute is the binary gender of the person.
+We use the same dataset splits as liu2021just.
+MetaShift.
+Given an image of an animal, the task is to predict whether it is a dog or cat [liang2022metashift].
+Following liang2022metashift, the attribute is the environment that the dog or cat is in.
+We use the same dataset splits as yang2023change.
+7 Supplementary Experiments
+7.1 Experimental Results
+Additional Dropping Ratios.
+We tabulate in Table 10 the results of additional dropping ratios $\{0.25,0.33,0.7\}$ as used by [minderer2024scaling].
+We observe a natural degradation in performance as the ratio increases with a sharp drop at 0.7.
+Likewise, in Figure 7, we find the image captions on COCO only begin to deviate significantly from $Base$ when dropping ratios are above 0.5.
+Power and Runtime.
+In Table 9, we provide the power (watt) and runtime (millisecond) that were used to calculate the energy (joule) consumed per sample in Table 2.
+Note that the runtime is first converted to seconds before taking its product with the power to produce the energy values.
+7.2 Ablation Studies
+Random Dropping of Tokens.
+In Table 11, we provide further analysis on the intra-image approach by randomly dropping tokens.
+This is similar to simoulin2024memory with the exclusion of any fine-tuning.
+First, unlike CelebA, we observe a noticeably degradation in worst-group accuracies on Waterbirds and MetaShift.
+We posit that this stems from the target label being spuriously correlated with the background (Waterbirds, MetaShift) and not gender (CelebA), which is distinctly separable from the foreground object as shown in Figure 3.
+Hence, random dropping is beneficial to the subgroup robustness of CelebA by reducing the gender features of the individuals.
+Second, we find performance (except ${\mathcal{T}}_{intra}^{0.7}$) to be slightly lower or equivalent on OpenImages and the remaining datasets, respectively.
+We attribute this to cases where the intra-image approach misidentifies the foreground object as irrelevant information in Figure 3 with the removal of the “airplane” (top image) and “building” (bottom image).
+However, in general, the variance of the individual patches remains effective for robustly compressing redundant information within the image.
+Fairness of Tokenization.
+Text tokenizers are known to induce unfairness between languages that raises compute costs, particularly for minority languages [petrov2024language, ali2023tokenizer].
+We seek to analyze if similar effects exist with VWTs.
+In Table 12, we show the breakdown in token length (including [CLS]) and accuracy (w.r.t $Base$) by subgroup.
+First, we observe a notable difference in compression between the subgroups of Waterbirds.
+With ${\mathcal{T}}_{inter}^{100}$, sequences might differ by up to 39 tokens as seen with subgroups 0 and 3.
+Smaller discrepancies are displayed on CelebA and MetaShift except for ${\mathcal{T}}_{inter}^{100}$ on the former.
+Second, we find compression to not affect all subgroups equally.
+Accuracy improves on certain subgroups and degrades on others.
+Stronger compression does not necessarily correlate with a larger change in performance.
+Sparsity of the Vocabulary.
+To better understand the utilization of the visual words, we plot the probability distribution of the matches in Figure 8.
+Regardless of the dataset, we find that certain visual words are matched more frequently than others, thus leading to a large skew in the distributions.
+Greater sparsity is also displayed by larger vocabularies as many visual words remain unused across the datasets.
+As such, the pruning of unmatched visual words may be applied to achieve a more efficient vocabulary size after the Pre-Process step of the inter-image approach.
+VWTs with Batching.
+In Figure 9, we seek to better understand the effectiveness of VWTs for offline inference where batch sizes are greater than 1.
+Using batches of $\{4,8,16,32\}$, we compare the energy (joule) consumed by the VWTs of ${\mathcal{T}}_{intra}^{0.5}$ and ${\mathcal{T}}_{inter}^{100}$ to $Base$ and $ToMe$.
+We find that as the batch size increases, both VWTs and $ToMe$ continue to generally improve energy efficiency across the datasets.
+This reduction in energy consumed is highest by ${\mathcal{T}}_{intra}^{0.5}$ followed by $ToMe$ and ${\mathcal{T}}_{inter}^{100}$.
+On Waterbirds, the inter-image approach does not result in a noticeable improvement or degradation in energy consumed.
+Dataset Subgroup Label Attribute Training Validation Test
+Waterbirds 0 0 (landbird) 0 (on land) 3498 467 2255
+1 0 (landbird) 1 (on water) 184 466 2255
+2 1 (waterbird) 0 (on land) 56 133 642
+3 1 (waterbird) 1 (on water) 1057 133 642
+CelebA 0 0 (non-blond) 0 (woman) 71629 8535 9767
+1 0 (non-blond) 1 (man) 66874 8276 7535
+2 1 (blond) 0 (woman) 22880 2874 2480
+3 1 (blond) 1 (man) 1387 182 180
+MetaShift 0 0 (dog) 0 (outdoor) 784 127 273
+1 0 (dog) 1 (indoor) 507 75 191
+2 1 (cat) 0 (outdoor) 196 33 65
+3 1 (cat) 1 (indoor) 789 114 345
+Table 8: Defined subgroups in Waterbirds, CelebA, and MetaShift.
+Dataset Model BaseBase $Q_{8}$ ToMeToMe ${\mathcal{T}}_{intra}^{0.5}$ ${\mathcal{T}}_{inter}^{100}$ ${\mathcal{T}}_{inter}^{1000}$ ${\mathcal{T}}_{inter}^{10000}$
+Waterbirds CLIP 123.99 73.89 91.64 110.93 102.65 107.93 117.36
+CelebA 126.53 74.47 93.07 114.37 96.31 102.12 115.46
+MetaShift 123.30 74.02 90.96 113.84 98.98 105.78 114.70
+OpenImages (Com.) 135.05 74.80 94.63 114.62 106.33 112.67 123.95
+OpenImages (Rare) 135.66 75.27 94.04 115.78 104.73 110.75 123.05
+COCO BLIP 191.74 81.68 147.23 169.77 144.71 156.25 171.84
+NoCaps 185.12 81.89 149.26 175.06 141.20 155.11 170.38
+(a) Power $\downarrow$
+Dataset Model BaseBase $Q_{8}$ ToMeToMe ${\mathcal{T}}_{intra}^{0.5}$ ${\mathcal{T}}_{inter}^{100}$ ${\mathcal{T}}_{inter}^{1000}$ ${\mathcal{T}}_{inter}^{10000}$
+Waterbirds CLIP 8.48 88.55 17.04 5.84 9.88 9.73 9.91
+CelebA 8.33 89.41 17.25 5.86 9.45 9.54 9.99
+MetaShift 8.21 89.48 17.23 6.01 9.99 9.87 9.75
+OpenImages (Com.) 8.83 92.23 17.44 5.97 9.64 9.69 10.23
+OpenImages (Rare) 8.60 86.00 17.04 6.19 9.69 10.02 9.67
+COCO BLIP 12.45 57.11 15.06 8.10 8.73 9.10 10.31
+NoCaps 12.44 57.61 15.08 8.04 9.49 9.28 10.51
+(b) Runtime $\downarrow$
+Table 9: Power (watt) and runtime (millisecond) per sample. The runtime is first converted to seconds before taking its product with the power to produce the energy values in Table 2.
+Dataset Model ${\mathcal{T}}_{intra}^{0.25}$ ${\mathcal{T}}_{intra}^{0.33}$ ${\mathcal{T}}_{intra}^{0.7}$
+Waterbirds CLIP 148 132 59
+CelebA
+MetaShift
+OpenImages (Com.)
+OpenImages (Rare)
+COCO BLIP 433 386 173
+NoCaps
+(a) Length
+Model Waterbirds CelebA MetaShift OpenImages
+Average $\uparrow$ Worst $\uparrow$ Average $\uparrow$ Worst $\uparrow$ Average $\uparrow$ Worst $\uparrow$ Common $\uparrow$ Rare $\uparrow$
+${\mathcal{T}}_{intra}^{0.25}$ 78.31 26.69 89.46 48.47 94.70 87.69 69.79 63.37
+${\mathcal{T}}_{intra}^{0.33}$ 77.72 27.78 89.31 48.83 94.62 87.18 69.11 62.75
+${\mathcal{T}}_{intra}^{0.7}$ 73.27 18.80 89.19 45.37 82.61 63.59 46.83 42.68
+(b) CLIP (image classification and subgroup robustness)
+Model COCO NoCaps
+Karpathy $\uparrow$ In-Domain $\uparrow$ Near-Domain $\uparrow$ Out-of-Domain $\uparrow$ Overall $\uparrow$
+BLEU@4 CIDEr CIDEr SPICE CIDEr SPICE CIDEr SPICE CIDEr SPICE
+${\mathcal{T}}_{intra}^{0.25}$ 33.92 107.47 103.59 14.65 98.39 14.07 94.47 13.59 98.34 14.06
+${\mathcal{T}}_{intra}^{0.33}$ 33.59 106.25 103.19 14.65 97.40 13.95 92.66 13.41 97.27 13.95
+${\mathcal{T}}_{intra}^{0.7}$ 29.64 93.20 89.89 13.53 85.04 12.82 80.81 12.31 84.88 12.82
+(c) BLIP (image captioning)
+Table 10: Token length (including [CLS]) and performance of the intra-image approach with varying dropping ratios. Performance degrades naturally as the ratio increases before falling sharply at 0.7.
+Figure 7: Visualization of image captions on COCO by the intra-image approach. The generated captions to not deviate significantly from those of BaseBase with dropping ratios of up to 0.5.
+Model Waterbirds CelebA MetaShift OpenImages
+Average $\uparrow$ Worst $\uparrow$ Average $\uparrow$ Worst $\uparrow$ Average $\uparrow$ Worst $\uparrow$ Common $\uparrow$ Rare $\uparrow$
+${\mathcal{T}}_{intra}^{0.25}$ 76.90 21.50 89.88 48.75 94.39 85.13 70.00 63.38
+${\mathcal{T}}_{intra}^{0.33}$ 75.72 20.51 90.10 48.89 93.44 84.62 69.66 63.06
+${\mathcal{T}}_{intra}^{0.5}$ 72.38 19.52 90.56 51.85 91.61 81.93 67.59 61.61
+${\mathcal{T}}_{intra}^{0.7}$ 66.78 11.58 90.24 61.48 84.29 69.73 52.72 49.86
+(a) CLIP (image classification and subgroup robustness)
+Model COCO NoCaps
+Karpathy $\uparrow$ In-Domain $\uparrow$ Near-Domain $\uparrow$ Out-of-Domain $\uparrow$ Overall $\uparrow$
+BLEU@4 CIDEr CIDEr SPICE CIDEr SPICE CIDEr SPICE CIDEr SPICE
+${\mathcal{T}}_{intra}^{0.25}$ 33.89 106.59 103.70 14.39 98.02 14.01 93.79 13.24 97.98 13.92
+${\mathcal{T}}_{intra}^{0.33}$ 33.78 106.57 102.44 14.35 98.15 13.98 93.04 13.21 97.73 13.88
+${\mathcal{T}}_{intra}^{0.5}$ 33.32 104.65 100.70 14.06 95.11 13.72 92.25 13.22 95.33 13.67
+${\mathcal{T}}_{intra}^{0.7}$ 31.15 97.18 95.02 13.57 86.83 13.01 86.43 12.52 87.93 13.00
+(b) BLIP (image captioning)
+Table 11: Token length (including [CLS]) and performance of the intra-image approach with random dropping.
+Unlike CelebA, the subgroup robustness degrades noticeably on Waterbirds and MetaShift.
+Performance is also slightly lower on OpenImages but equivalent on the remaining datasets except for ${\mathcal{T}}_{intra}^{0.7}$.
+Model Subgroup Waterbirds CelebA MetaShift
+Length $\Delta$ Accuracy Length $\Delta$ Accuracy Length $\Delta$ Accuracy
+BaseBase 0 197 98.89 197 95.63 197 98.54
+1 82.45 96.23 93.19
+2 21.86 48.67 87.69
+3 54.73 50.00 95.36
+${\mathcal{T}}_{inter}^{100}$ 0 144 -0.67 85 -3.25 118 -0.74
+1 106 -3.82 88 -0.13 110 -6.18
+2 140 19.00 98 32.62 119 -2.92
+3 105 6.17 99 6.30 100 -9.02
+${\mathcal{T}}_{inter}^{1000}$ 0 160 -0.40 119 -0.46 138 -0.25
+1 129 -0.22 117 0.74 138 -2.06
+2 160 9.03 124 18.92 143 -1.17
+3 129 2.18 124 -4.07 131 -2.13
+${\mathcal{T}}_{inter}^{10000}$ 0 180 -0.09 158 0.25 163 -0.37
+1 156 1.34 151 0.66 167 0.19
+2 181 4.75 158 2.96 169 -1.75
+3 158 1.71 156 -6.30 163 -0.81
+Table 12: Distribution of token length (including [CLS]) and accuracy (w.r.t.
+BaseBase) by subgroup.
+${\mathcal{T}}_{inter}^{{\mathcal{V}}}$ of varying vocabulary sizes are shown.
+Like text tokenizers, VWTs may induce unequal token lengths as seen with ${\mathcal{T}}_{inter}^{100}$ on Waterbirds.
+Performance is also affected unequally as a stronger sequence compression does not correlate with a greater improvement or degradation in accuracy.
+(a) Waterbirds
+(b) CelebA
+(c) MetaShift
+Figure 8: Probability distribution of the matched visual words.
+${\mathcal{T}}_{inter}^{{\mathcal{V}}}$ of varying vocabulary sizes are shown.
+The probability distribution exhibits a large skew irrespective of the dataset as certain visual words are matched more frequently than others.
+Larger vocabularies display greater sparsity as the many visual words that remain unmatched may be pruned for a more efficient vocabulary size.
+[…]